### v1

# Hybrid Deepfake Image Detection: A Comprehensive Dataset-Driven Approach Integrating Convolutional and Attention Mechanisms with Frequency Domain Features

Abstract
Effective deepfake detection tools are becoming increasingly essential over the last few years due to the growing usage of deepfakes in unethical practices. There exists a diverse range of deepfake generation techniques, which makes it challenging to develop an accurate universal detection mechanism. The 2025 Signal Processing Cup (DFWild-Cup competition) provided a diverse dataset of deepfake images, which are generated from multiple deepfake image generators, for training machine learning model(s) to emphasize the generalization of deepfake detection. To this end, we proposed an ensemble-based approach that employs three different neural network architectures: a ResNet-34-based architecture, a data-efficient image transformer (DeiT), and an XceptionNet with Wavelet Transform to capture both local and global features of deepfakes. We visualize the specific regions that these models focus for classification using Grad-CAM, and empirically demonstrate the effectiveness of these models in grouping real and fake images into cohesive clusters using t-SNE plots. Individually, the ResNet-34 architecture has achieved 88.9% accuracy, whereas the Xception network and the DeiT architecture have achieved 87.76% and 89.32% accuracy, respectively. With these networks, our weighted ensemble model achieves an excellent accuracy of 93.23% on the validation dataset of the SP Cup 2025 competition. Finally, the confusion matrix and an Area Under the ROC curve of 97.44% further confirm the stability of our proposed method.
keywords:
Deepfake , ResNet-34 , Transformer , DeiT , Wavelet , XceptionNet
journal: Elsevier Image and Vision Computing
[…]
[inst1]organization=Department of Electrical and Electronic Engineering,addressline=BUET, city=Dhaka, postcode=1205, country=Bangladesh
[…]
[inst2]organization=School of Data and Sciences,addressline=BRAC University, city=Dhaka, postcode=1212, country=Bangladesh
1 Introduction
Deepfakes are media that are edited or created by generative Artificial Intelligence (AI), which do not exist or have not occurred in reality. Although deepfakes have existed for a long time, it took experts, and often entire studios, to generate convincing fake media. Due to the surge of machine learning and generative AI tools, the generation of deepfakes has never been easier and more accessible. Such ease of access has raised serious societal concerns since these are being used to manipulate public opinion, commit fraud, create non-consensual pornography, and attack individuals through defamation or identity threats [1]. Deepfakes are also being used to gain malicious political advantage worldwide [2]. Although deepfake generation tools are widely accessible, detection tools are not nearly as effective and accurate. More specifically, the existing detection methods only work well when the test image resembles the features of the images of the training dataset of the detector. That is, the major issues in the existing detection approaches are the lack of diversity in training datasets, and the inability to be up to date with new fake image generation techniques [3].
To address the research gap, the DFWild-Cup competition called for novel approaches that can achieve accurate deepfake detection and provided a diverse dataset (consisting of samples from eight different standard datasets) to help train a generalized deepfake detection algorithm. Such a diverse dataset poses unique challenges to formulating accurate image detection methods. In our proposed solution for deepfake image detection, we have employed a weighted ensemble approach, which utilized three different neural network architectures with different strengths: i) a ResNet-34-based architecture with squeeze and excitation block, ii) a data-efficient image transformer (DeiT) model, and iii) an Xception network. More specifically, we consolidate the outputs from these architectures and employ a weighted ensemble mechanism to get the final result. We note that the ResNet-34-based architecture, due to its residual connections, allows learning complex hierarchical features with deeper networks by addressing the problem of vanishing gradients [4]. Moreover, the squeeze and excitation (SE) block incorporated in the ResNet-34-based architecture provides channel-wise attention with very little computational expense [5]. Besides, the self-attention mechanism and the distillation token introduced in the DeiT model allow learning long-range dependencies across the entire image. More specifically, the ResNet model excels at capturing local spatial relationships of an image, while the transformer model can capture global features of an image through the self-attention mechanism [6]. Finally, the Xception model, combined with wavelet transform as a preprocessing technique, is effective in detecting fine-grained artifacts in deepfake images. We have utilized the ’Haar’ wavelets due to their notable ability to detect abrupt transitions and edges in the data [7]. A weighted ensemble mechanism of these three networks is used as the final solution, which has enabled us to achieve superior performance compared to the models operating alone.
Related Works. The advancement in deepfake research has led to the creation of nearly perfect image manipulations that are undetectable to the human eye. The combination of multiple deep learning architectures has been shown to perform well on deepfake detection. Gowda and Thillaiarasu [8] showed ensemble techniques of Convolutional Neural Network (CNN) models (ResNeXt and Xception) achieve higher accuracy on deepfake detection. Wodajo and Atnafu [9] combined VGG-16 for feature extraction and Vision Transformer (ViT) for classification, showing promising results on the DFDC dataset. Wolter et al. [10] proposed a wavelet-packet-based analysis for detecting GAN-generated images, highlighting spatial frequency differences in synthetic content and achieving competitive results using combined architectures of wavelets and CNNs. Ricker et al. [11] proposed the usage of frequency domain features extracted via DCT, and achieved high accuracy rates (97.7% for GAN and 73% for diffusion models), highlighting the utility of frequency-domain artifacts. Patel et al. developed an enhanced deep CNN (D-CNN) architecture for deepfake detection, achieving reasonable accuracy and high generalization capability [12]. Patch-Forensics, a technique analyzing smaller patches to detect local image artifacts, achieves 100% average precision on StyleGAN-generated images using an Xception Block 2 classifier [13]. However, Abdullah et al. [14] highlighted two major limitations of the existing networks arguing that the state-of-the-art models often lack control over content and image quality, making them ineffective at detecting high-quality deepfakes if trained on low-quality fake images, and they also lack adversarial evaluation, where an attacker can exploit knowledge of the defense. Works on adversarial perturbation attacks to fool deep neural network (DNN) based detectors are a challenge in deepfake detection [15, 16]. Training on diverse and adversarially perturbed datasets can increase the robustness of a detection architecture [17]. Lin et al. proposed a CNN-based method for deepfake detection that combines multi-scale convolution with a vision transformer [18]. Their approach features a multi-scale module incorporating dilation convolution and depth-wise separable convolution. This design aims to capture more facial details and identify tampering artifacts at various scales [18]. Note that the ViT has gained significant popularity in recent years for deepfake image detection [19, 20].
Researchers are not only relying on complex network architectures but also exploring other approaches, such as the Frequency Enhanced Self-Blending Images (FSBI) method for deepfake detection, which utilizes the Discrete Wavelet Transform (DWT) [21]. Zhu et al. introduced a learnable Image Decomposition Module (IDM) that uses a recomposition operation to highlight illumination inconsistencies [22]. It features a multi-level enhancement technique for effective feature recomposition and is trained in the logarithmic domain, with an extensible design [22]. To address challenges with low-quality datasets and inadequate detection performance, Cheng et al. have developed MSIDSnet [23]. This framework utilizes a multi-scale fusion (MSF) module to capture forged facial features and an interactive dual-stream (IDS) module to enhance feature integration across frequency and spatial domains [23]. Another recent detection model by Byeon et al. uses a spatial-frequency joint dual-stream convolutional neural network with learnable frequency-domain filtering kernels [24].
2 Proposed Methodology: Weighted Ensemble of Neural Network Models
2.1 Description of the Dataset
As mentioned before, the DFWild-Cup competition has provided a diverse dataset. The dataset contains images from eight publicly available standard datasets designed for the DeepfakeBench evaluation [25]: Celeb-DF-v1 [26], Celeb-DF-v2 [26], FaceForensics++ [27], DeepfakeDetection [28], FaceShifter [29], UADFV [30], Deepfake Detection Challenge Preview [31], and Deepfake Detection Challenge [32]. To ensure that the detector performance is not biased by specific pre-processing techniques, all datasets underwent uniform pre-processing. Additionally, any information that would identify the origin of the dataset has been eliminated, and the file names have been anonymized. For training, there are a total of 42,690 real images and 219,470 fake images. The validation dataset consists of 1,548 real images and 1,524 fake images. It is evident that there is a significant class imbalance present in the training data as the ratio of real and fake images is approximately 5:1.
Figure 1: Training procedure including class imbalance handling
2.2 Data Preprocessing
2.2.1 Class Balancing for Model Training
Since the number of fake images is approximately five times the number of real images, we divided the fake images ${}\{\mathcal{F}\}$ into five disjoint subsets $\{\mathcal{F}_{1},\ldots,\mathcal{F}_{5}\}$. Each of the three aforementioned models, denoted as $M_{1},M_{2},M_{3}$, was trained in five stages with these disjoint subsets of fake images. At each stage $i$, the model is trained using the entire real image set $\mathcal{R}$ and one subset of fake images $\mathcal{F}_{i}$, such that the training dataset of stage $i$ is given by:
$\mathcal{D}_{i}=\mathcal{R}\cup\mathcal{F}_{i},\quad i\in\{1,2,3,4,5\}.$ (1)
We emphasize that the same set of real images ${}\{\mathcal{R}\}$ are treated as one class during each stage. The training of each model is initialized with pre-trained weights $\theta_{0}$, obtained from an ImageNet [33]-trained version of the corresponding architecture. The model parameters at each stage are updated iteratively using the dataset $\mathcal{D}_{i}$, following the optimization process:
$\theta_{i}=\arg\min_{\theta}\sum_{(x,y)\in\mathcal{D}_{i}}\mathcal{L}(M(x;\theta),y),$ (2)
where $\mathcal{L}$ represents the chosen loss function. After training on $\mathcal{D}_{i}$, the final weights $\theta_{i}$ from each stage are used to warm-start the next stage:
$\theta_{i+1}\leftarrow\theta_{i},\quad\text{for }i\in\{1,2,3,4\}.$ (3)
This process continues until the model has been trained on all subsets of fake images, ensuring that the model learns from diverse variations of the fake images while preserving knowledge of real images. Figure 1 illustrates the sequential learning procedure.
2.2.2 Data Augmentation
As mentioned before, we utilized the same set of real images across five training stages to address the class imbalance. However, this evidently increases the chance of overfitting. To mitigate overfitting and enhance the model’s generalization performance, we implemented four data augmentation techniques: random horizontal flips, random rotations, color jittering, and random resized cropping. The random horizontal flip method flips the training images along their vertical axis with a probability of 50%. The Random rotation method augments images by randomly rotating them within $0^{o}-10^{o}$. The color jitter method randomly alters the brightness, contrast, saturation, and hue of images within a specified range. To this end, a range of 0.2 for brightness, contrast, saturation, and hue provided the best results in our experiments. Finally, the random resized crop was used as a data augmentation technique that applies both random cropping and resizing to create varied versions of training images. In our case, it produced $224\times 224$ images, while randomly scaling the crop area between 80% and 120% of the original image size. These augmentations enhanced the diversity of the training dataset, reduced overfitting, and improved the model’s ability to generalize across a wider range of deepfake detection scenarios. It should be noted that such augmentations were used only in training the ResNet-34 and the DeiT architectures. For the XceptionNet, we employed wavelet transform-based feature extraction from the raw images, and data augmentation techniques would result in incorrect feature representation in the spectral domain.
2.2.3 Data Resizing and Normalization
All three models were pre-trained using the ImageNet dataset. Since the requirements of input image dimensions are different on the ResNet-34 architecture and the Xception network, we have resized the input images accordingly to the specific dimensions for respective architectures. Additionally, the training images were normalized using the mean and standard deviation of the ImageNet dataset [33].
2.3 Wavelet Feature Extraction for XceptionNet
The wavelet transform is shown to be effective for image feature extraction as it decomposes images into multi-resolution frequency components [34]. The ability to decompose an image into different frequency sub-bands while preserving spatial relationships makes it particularly effective at identifying edges, textures, and subtle patterns of an image [35]. Using 2D Wavelet transform, an image can be decomposed into approximate and detail coefficients [7]. As such, we performed a single-level 2D Discrete Wavelet Transform (2D-DWT) using Haar wavelets to produce the approximate and detailed coefficients of the images and used the computed coefficients as features.
Figure 2: Row and Column-wise filtering and downsampling to generate approximation and detail images through Haar 2D-DWT
The 2D-DWT is applied to an image $I(x,y)$ to decompose it into frequency sub-bands while preserving spatial information. Given an image $I\in\mathbb{R}^{M\times N}$, a single-level 2D-DWT using Haar wavelets generates four coefficient matrices: the approximate coefficients and three sets of detail coefficients (horizontal, vertical, and diagonal). The wavelet decomposition is performed using separable filtering along rows and columns. The Haar wavelet uses two filters:
  • 1.
    Low-pass filter ($\phi$): Coefficients $\left[\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right]$
  • 2.
    High-pass filter ($\psi$): Coefficients $\left[-\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right]$
As shown in Figure 2, the sub-bands are computed as follows:
  1. 1.
    Row-wise filtering and downsampling:
    $L_{\text{row}}$ $=\left(I\ast\phi\right)\downarrow_{2,\text{cols}}$
    $H_{\text{row}}$ $=\left(I\ast\psi\right)\downarrow_{2,\text{cols}}$
  2. 2.
    Column-wise filtering and downsampling:
[…]
where $\ast$ indicates convolution; $\downarrow_{2,\text{cols}}$, and $\downarrow_{2,\text{rows}}$ indicate downsampling the columns and rows by 2, respectively; $A$ denotes low-frequency approximation coefficients; and $H,V,D$ denote high-frequency horizontal, vertical, and diagonal detail coefficients, respectively. For an RGB image $I_{r},I_{g},I_{b}$ consisting of three color channels, we apply the above transformation independently to each color channel:
$A_{c},H_{c},V_{c},D_{c}=\text{DWT}(I_{c}),\quad c\in\{r,g,b\},$ (4)
where $A_{c},H_{c},V_{c},D_{c}$ are the approximate and detail coefficients for each color channel $c$. The final feature representation is formed by concatenating the coefficients across all three channels:
$F=\text{concat}((A_{r},A_{g},A_{b}),(H_{r},H_{g},H_{b}),(V_{r},V_{g},V_{b}),(D_{r},D_{g},D_{b})).$ (5)
To prepare the wavelet-domain features for classification, the coefficients are scaled to an 8-bit integer range and resized to match the input dimensions of the XceptionNet model. In Figure 3, we showed two such feature images with the approximate and detailed coefficients of two sample images from the dataset.
Figure 3: Real and Fake feature images after wavelet transform. The top left, top right, bottom left, and bottom right portions indicate the approximate, horizontal detail, vertical detail, and diagonal detail coefficients of the original image, respectively. Different colormaps are applied to the horizontal, vertical, and diagonal detail coefficients, and their contrasts are increased for better visibility.
2.4 Details of the Model Architectures
Recall that our proposed solution utilizes three neural network models for a weighted ensemble approach for our final prediction. We chose each of these models to focus on different aspects of a deepfake image so that the ensemble approach provides better accuracy than individual models operating alone. The details of the models in our proposed solution are presented below.
2.4.1 ResNet-34 with SE Block
We implemented a ResNet-34 model with SE block for the detection of deepfake images. Note that, we investigated with different Convolutional Neural Network (CNN)-based architectures – shallow CNNs exhibited high bias, whereas deep CNNs caused vanishing gradients. The residual mapping of the ResNet-34 architecture addresses the issue of vanishing gradients, while the capacity of the ResNet-34 architecture is sufficient for capturing the intricate features for deepfake classification.
Figure 4: Building block of residual learning
Note that residual learning works by reformulating the target function $\text{H}(x)$ as $\text{F}(x)=\text{H}(x)-x$, where $x$ is the input [4]. Instead of directly learning $\text{H}(x)$, the network learns the residual $\text{F}(x)$. The final output is computed as $\text{F}(x)+x$, combining the learned residual with the input via a skip connection. This mechanism allows each layer to focus on learning the residual adjustments needed to improve upon the input. The building block of residual mapping is shown in Figure 4. Our ResNet-34-based architecture is implemented with two simple principles following the traditional structure of this model: i) When maintaining the same output feature map size, each layer will consistently have the same number of filters. ii) Additionally, when the feature map size is halved, the number of filters is effectively doubled. Based on these principles, shortcut connections are implemented on a generic CNN architecture to obtain the desired residual network.
Figure 5: ResNet-34 architecture with Squeeze and Excitation
Squeeze and Excitation (SE) Block. The SE block is introduced as a lightweight attention mechanism to adaptively recalibrate each feature channel’s importance [5]. The SE block is placed after the final sequential layer (layer 4) of the ResNet-34 model to enhance feature representation by modeling channel-wise dependencies. It is responsible for two main operations:
  • 1.
    Squeeze: Global spatial information is aggregated using global average pooling to generate channel-wise descriptors. Given an input feature map $X\in\mathbb{R}^{H\times W\times C}$, the squeezed feature vector $z\in\mathbb{R}^{C}$ is obtained by squeezing function $F_{sq}(.)$ as:
    $z_{c}=\frac{1}{HW}\sum_{i=1}^{H}\sum_{j=1}^{W}X_{c}(i,j),\quad c=1,2,...,C,$ (6)
    where $X_{c}(i,j)$ represents the activation at spatial location $(i,j)$ in channel $c$.
  • 2.
    Excitation: The descriptors generated from the Squeeze block are passed through a lightweight two-layer fully connected network with a bottleneck structure to model non-linear channel independencies. The squeezed feature vector is passed through a gating mechanism with two fully connected layers and a non-linearity (typically ReLU and Sigmoid) to learn channel-wise dependencies:
    $s=\sigma(W_{2}\delta(W_{1}z)),$ (7)
    where $W_{1}\in\mathbb{R}^{\frac{C}{r}\times C}$ and $W_{2}\in\mathbb{R}^{C\times\frac{C}{r}}$ are the weights of the two fully connected layers, $\delta(\cdot)$ represents the ReLU activation function, and $\sigma(\cdot)$ is the Sigmoid function that outputs a channel-wise scaling factor $s\in\mathbb{R}^{C}$. The final output is obtained by rescaling the input feature maps $X$ using the channel-wise attention weights:
    $\tilde{X}_{c}=s_{c}X_{c},\quad c=1,2,...,C$ (8)
    This step selectively enhances informative channels while suppressing less important ones. We denoted this operation as a function $F_{ex}(.)$, as depicted in Fig. 5.
The details of the ResNet-34 architecture including the output shapes of each layer and the number of parameters are shown in Table 1.
Table 1: Details of the ResNet-34 Architecture with SE Block
Layer (type) Output Shape Param #
Conv2d, BatchNorm2d, ReLU [64, 112, 112] 9,536
MaxPool2d [64, 56, 56] 0
BasicBlock-1 $\times$ 3 [64, 56, 56] 221,952
   —Conv2d [64, 56, 56] 36,864
   —BatchNorm2d + ReLU [64, 56, 56] 128
   —Conv2d [64, 56, 56] 36,864
   —BatchNorm2d + ReLU [64, 56, 56] 128
BasicBlock-2 $\times$ 4 [128, 28, 28] 1,116,416
BasicBlock-3 $\times$ 6 [256, 14, 14] 6,822,400
BasicBlock-4 $\times$ 3 [512, 7, 7] 13,114,368
SEBlock (Global Avg Pool + FC) [512, 7, 7] 33,312
AdaptiveAvgPool2d [512, 1, 1] 0
Linear [2] 1,026
Total Parameters 21,319,010
Trainable Parameters 8,204,642
Non-Trainable Parameters 13,114,368
2.4.2 Data Efficient Image Transformer (DeiT) Model
Unlike conventional Recurrent Neural Networks (RNNs) that process sequential data step by step, transformers process all inputs simultaneously, relying on a self-attention mechanism. This simultaneous processing allows transformers to learn long-range dependencies and global contexts of data [6]. It has been shown that the data-efficient image transformer, which is based on the backbone of the vision transformer, works well for cases where the training data is limited [36]. Here, an image is divided into a grid of smaller patches as a sequence of input tokens ($N$ number of $16\times 16$ patches), and each patch is associated with a vector through linear embedding. After adding positional encoding for each patch, the vectors are fed into the transformer encoder, where the self-attention mechanism helps to learn the relationship among all the patches [37]. Note that the attention mechanism is based on a trainable associative memory of (key, value) pairs. A query vector $\mathbf{Q}$ is matched against key vectors $\{\mathbf{K}\}$ using dot products, scaled by $\sqrt{d}$ and normalized via a softmax function to produce weights. These weights are used to compute a weighted sum of value vectors $\{\mathbf{V}\}$, producing an output matrix as:
$\text{attention}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\mathbf{V}.$
In self-attention, the queries, keys, and values are derived from the same input sequence via linear transformations [38]. Multi-head self-attention (MSA) applies $h$ separate attention functions (heads) in parallel. Each head provides a sequence of outputs, which are concatenated and projected back to the original dimension. This allows the model to attend to different parts of the input sequence simultaneously.
Figure 6: Data Efficient Image Transformer (DeiT) architecture
Note that CNN models primarily capture local contexts and eventually may capture global contexts through hierarchical learning. To this end, DeiT is an effective alternative to traditional CNNs in learning global features [6]. It uses knowledge distillation techniques and optimization strategies to improve generalization. DeiT learns from a pre-trained CNN model (that is, the RegNetY-16GF model [39]) and uses a distillation token during training, allowing it to learn faster and more efficiently with less data [36].
The DeiT architecture introduces a new hard distillation mechanism where they treat the hard decisions of the teacher model $y_{t}$ as true labels along with the actual true labels, $y$. If $Z_{t}$ are the logits of the teacher model and $Z_{s}$ are the logits of the student model, then the hard decision of the teacher model is defined as
$y_{t}=\text{argmax}_{c}(Z_{t}(c))$ (9)
Here $c$ is the class index. Using this prediction, the final loss function for hard distillation is treated as
$L^{\text{hard dist}}_{\text{global}}=\frac{1}{2}L_{CE}(\text{softmax}(Z_{s}),y)+\frac{1}{2}L_{CE}(\text{softmax}(Z_{s}),y_{t})$ (10)
The first term of the equation can be thought of as $L_{\text{cross-entropy}}$ and the second term as $L_{\text{teacher}}$ depicted in Figure 6
As the competition dataset lacks enough samples to generalize a ViT, we chose DeiT to capture global contexts effectively, despite it being computationally costly. There exist different DeiT model variations based on the number of heads and embedding resolutions. In our proposed solution, we have implemented the DeiT-B model, which uses the ViT-B model as its backbone. It has 768 embedding dimensions with 12 heads containing a total of 86 million parameters [36]. The DeiT model architecture is shown in Figure 6 and the number of parameters and the output shapes of each layer are shown in Table 2.
Table 2: Details of the Data Efficient Image Transformer (DeiT) Model
Layer (type) Output Shape Param #
Conv2d [764, 14, 14] 590,592
Patch Embed [196, 768] 0
Dropout [197, 768] 0
Identity-1 [197, 768] 0
LayerNorm $\times$ 25 [197, 768] 38,400
Linear-1 $\times$ 12 [197, 2304] 21,261,312
Identity-2 [12, 197, 64] 0
Linear-2 $\times$ 12 [197, 768] 35,407,872
Attention $\times$ 12 [197, 768] 0
Linear-3 $\times$ 12 [197, 3072] 28,348,416
GELU $\times$ 12 [197, 3072] 0
Dropout [197, 3072] 0
Mlp $\times$ 12 [197, 768] 0
Identity and Dropout [768] 0
Linear-4 [2] 1,538
Vision Transformer [2] 0
Total Parameters 85,648,130
Trainable Parameters 85,648,130
Non-Trainable Parameters 0
2.4.3 XceptionNet Model
XceptionNet is based on the concept of depthwise separable convolutions, which enhance computational efficiency while maintaining accuracy. In depthwise convolution, a single convolutional filter is applied to each input channel. Following this, pointwise convolution is used to combine the outputs of the depthwise convolution into a linear form [40]. XceptionNet features a fully convolutional architecture for simplicity and efficiency, leverages residual connections to improve gradient flow, and employs deep feature extraction.
(a) Entry Flow
(b) Middle Flow
(c) Exit Flow
Figure 7: Details of XceptionNet architecture
As shown in Figure 7, our implemented XceptionNet has 36 convolutional layers, structured into 14 modules with residual connections, except for the first and last modules [40]. It has three primary sections—Entry Flow, Middle Flow, and Exit Flow. The Entry Flow layer consists of convolutional layers followed by max pooling and is responsible for reducing spatial dimensions and extracting low-level features. The Middle Flow layer contains multiple depthwise convolution blocks with residual connection, focusing on high-level features. The Exit Flow layer includes more depthwise separable convolutions and reduces spatial dimensions further. We modified the fully connected layer from the original architecture to match our binary classification problem. The summary of the output shapes and number of parameters of each block and layer is shown in Table 3.
Table 3: Details of the XceptionNet Model
Layer (type) Output Shape Param #
Conv2d [32, 149, 149] 864
BatchNorm2d + ReLU [32, 149, 149] 64
Conv2d [64, 147, 147] 18,432
BatchNorm2d + ReLU [64, 147, 147] 128
Block-1 [128, 74, 74] 35,264
   —SeparableConv2d [128, 147, 147] 8,768
   —BatchNorm2d + ReLU [128, 147, 147] 256
   —SeparableConv2d [128, 147, 147] 17,536
   —BatchNorm2d [128, 147, 147] 256
   —MaxPool2d [128, 74, 74] 0
   —Conv2d [128, 74, 74] 8,192
   —BatchNorm2d [128, 74, 74] 256
Block-2 [256, 37, 37] 136,064
Block-3 [728, 19, 19] 915,944
Block-4 $\times$ 8 [728, 19, 19] 12,911,808
Block-5 [1024, 10, 10] 2,039.584
SeparableConv2d [1536, 10, 10] 1,582,080
BatchNorm2d + ReLU [1536, 10, 10] 3072
SeparableConv2d [2048, 10, 10] 3,159,552
BatchNorm2d + ReLU [2048, 10, 10] 4096
SelectAdaptivePool2d [2048] 0
Linear [2] 4098
Total Parameters 20,811,050
Trainable Parameters 20,811,050
Non-Trainable Parameters 0
2.4.4 Weighted Ensemble Mechanism
Finally, we implemented a weighted ensemble mechanism to combine the predictions of the aforementioned three models. The associated weights for the ResNet-34 with SE, DeiT, and Xception with wavelet transform are 0.36, 0.45, and 0.19, respectively, based on our parameter search. The final architecture is shown in Figure 8, and the total number of trainable and non-trainable parameters in the final architecture is shown in Table 4.
Figure 8: The proposed weighted ensemble mechanism
Table 4: Number of parameters in the final weighted ensemble mechanism
Model Total Parameters Trainable Non-trainable
ResNet 21,319,010 8,204,642 13,114,368
DeiT 85,648,130 85,648,130 0
Xception 20,811,050 20,811,050 0
Total 127,778,990 114,664,622 13,114,368
3 Experimental Results
Figure 9: (a) Confusion matrix and (b) ROC curve for validation set
In this section, we demonstrate the performance of our proposed solution on the competition dataset. We note that investigation and experimentation with different architectures helped us make informed decisions to reach our proposed solution. More specifically, we started experimenting with simpler CNN models. As mentioned before, the performance of CNN models, such as the VGG-19, indicated insufficient capacity and indicated the necessity of deeper networks. Additionally, residual networks, such as the ResNet-50, raised overfitting issues. These results motivated us to choose a ResNet-34-based architecture to strike a balance between high-bias and high-variance regimes. We investigated the performance of the convolutional block attention module (CBAM) and SE and opted for SE block based on the generalization performance on the given dataset. Lastly, we investigated the effect of different handcrafted feature extraction methods on the performance and opted for wavelet feature extraction for our XceptionNet model.
(a) SEResNet (Untrained)
(b) DeiT (Untrained)
(c) Xception (Untrained)
(d) SEResNet (Trained)
(e) DeiT (Trained)
(f) Xception (Trained)
Figure 10: t-SNE plot of three models before and after training for randomly selected 4000 real and fake images using perplexity = 40
In essence, gradually increasing the complexity of networks helped us understand the shortcomings of different architectures for our specified task and led us to choose the model architectures based on their generalization capability with a relatively small dataset. To further reduce the variance, we employed a weighted ensemble of ResNet-34 with SE, DeiT, and XceptionNet to generate the final results. We present a comparative analysis of the results of different model variations and configurations that guided us to our final design decisions in Table. 5. We argue that these results serve as empirical evidence of our choice of neural network architectures, and the weighted ensemble approach.
Table 5: Comparative Analysis of Different Model Variations
Model Validation Acc.
Xception-Net [40] 61.48%
VGG-19 [41] 62.25%
ResNet-50 [4] 69.00%
DCT with ResNet-34 71.48%
ResNet-34 [4] 80.32%
ResNet-34 with CBAM 85.19%
Xception-Net with Wavelet Transform 87.76%
ResNet-34 with SE 88.90%
DeiT [36] 89.32%
Maj. Voting (ResNet, DeiT, Xception) 91.34%
Weighted Ensem. (ResNet, DeiT) 91.80%
Weighted Ensem. (ResNet, DeiT & Xception) (Proposed) 93.23%
Hyperparameters. For the training of our models, we have used the ADAM optimizer. We have used accuracy, precision, recall, and F1 score as our evaluation metrics. Since the deepfake detection problem is a binary classification problem, we used binary categorical cross-entropy as our loss function. Using the validation set to observe the accuracy of the proposed architecture, each image took approximately 0.03515625 sec on a Tesla P100 GPU. The hyperparameters used for training are presented in Table 6.
Table 6: Hyperparameters for model training
Hyperparameters Details
Batch size 32
Number of Epochs
   - ResNet-34 with SE 10
   - DeiT 5
   - XceptionNet 10
Learning rate
   - Learning rate scheduler ReducedLROnPlateau
   - Initial learning rate 0.0001
   - Reduction factor 0.1
   - Patience 3
We have utilized the NumPy and Pandas libraries for preprocessing and data manipulation tasks, and PyTorch [42] for model implementation, training, and validation. More specifically, we used torch for core functionality, torch.nn for neural networks, torch.optim for optimization, and utilities like torch.utils.data.Dataset, DataLoader, and random_split. From TorchVision, we used transforms for image transformations and models for pre-trained architectures. For performance metrics, we applied Scikit-learn (precision_score, recall_score, and f1_score). Progress tracking was managed with TQDM, and TorchSummary provided model summaries. Image processing utilities include OpenCV, while PyWavelets was utilized for wavelet transformations. Additionally, Timm was used to access advanced pre-trained models, and the Random module was applied for generating randomness.
(a) SEResNet
(b) DeiT
(c) Xception
(d) SEResNet
(e) DeiT
(f) Xception
Figure 11: Importance map shown by Grad-CAM for different models
(a) SEResNet
(b) DeiT
(c) Xception
Figure 12: Grad-CAM view of a few failed classifications of our proposed model
Result Analysis. As mentioned before, we evaluated our weighted ensemble model for multiple evaluation metrics, including accuracy, precision, recall, F1-score, area under receiver operating characteristic (AU-ROC) curve, equal error rate, and average precision. While calculating these metrics, fake images were considered as class 0 and real images as class 1. The performance of our model on these metrics is presented in Table 7 and the confusion matrix and the ROC curve are shown in Figure 9. It is evident from the evaluation indices that our proposed weighted ensemble architecture performs very well on the competition dataset, which underlines the generalization capability of the proposed approach.
Table 7: Performance of the Final Ensemble Architecture combining ResNet-34, DeiT, and XceptionNet
Evaluation Metrics Performance
Accuracy 93.23%
Precision 0.9515
Recall 0.9121
F1 score 0.9314
Area under ROC 0.9769
Equal Error Rate 6.33%
Average Precision 0.9744
To investigate the performance of our ensemble model, we intend to visualize i) the regions of the images that the three models focus for classification, and ii) how the three models distinguish real and fake images before and after training. To that end, we note that the Grad-CAM visualizations show how different models focus on different facial regions of the same image for deepfake detection. In Figure 11, it can be seen that the ResNet-34 and Xception models focus on localized facial features, whereas DeiT displays diffuse attention patterns indicating a focus on global features. Each of the models focuses on different parts of the image, capturing unique and complementary information. When these diverse insights are combined into an ensemble, the overall prediction becomes more accurate, robust, and better at generalizing to new data. Additionally, we present the Grad-CAM view of some of the wrong predictions by the three models in Figure 12, showing the focus of irrelevant regions, leading to mis-classifications.
Next, we note that high-dimensional data can be visualized by projecting onto a two-dimensional space using the non-linear dimensionality reduction technique known as t-SNE (t-Distributed Stochastic Neighbor Embedding) [43]. By maintaining the local neighborhoods of data points, it shows the differences and similarities of data points in the learned feature space. As such, the t-SNE visualizations (Fig. 10) offer valuable insights into the effectiveness of our models. For both real and fake images, we observe a single cohesive cluster for all three models. Notably, the cluster of fake images generated by two of our models, SEResNet and XceptionNet, is compact, despite these images being produced using various deepfake generation methods. The plots show that the training has effectively enhanced the ability to distinguish between real and fake image features for all three models. This, in-turn, highlights the generalization capability of our proposed ensemble model. After training, all three backbone models—SEResNet, DeiT, and XceptionNet—produced well-separated embeddings for real and fake images, as illustrated in Fig. 10(d), 10(e), and 10(f).
4 Conclusion
In this work, we have proposed a weighted ensemble approach that takes advantage of the strengths of a ResNet-34-based model, a DeiT model, and an XceptionNet model for the accurate classification of deepfake images. The ResNet-34 model enabled extracting local features from an image, while the DeiT model specialized in capturing global features. Additionally, XceptionNet, combined with wavelet transform, effectively identified frequency-dependent artifacts. We emphasize that detecting deepfake images that are generated with a range of generative AI approaches is a challenging task. Our proposed model architectures captured the intricate properties of the input images to successfully distinguish between real and fake images and produced excellent results. Through t-SNE plots, we demonstrated that our model is capable of separating real and fake images into separable clusters. Additionally, we included a Grad-CAM visualizations to map the focus-regions of images to illustrate how our models analyze the images and which portions contribute to distinguishing between real and fake images. An interesting future work could be the incorporation of lightweight architectures and reduce the computational overhead, as well as an increase in the model’s adaptability to newer and sophisticated deepfake generation methods.

### v2

@@
-Hybrid Deepfake Image Detection: A Comprehensive Dataset-Driven Approach Integrating Convolutional and Attention Mechanisms with Frequency Domain Features
+CAE-Net: Generalized Deepfake Image Detection using Convolution and Attention Mechanisms with Spatial and Frequency Domain Features
 Abstract
-Effective deepfake detection tools are becoming increasingly essential over the last few years due to the growing usage of deepfakes in unethical practices.
-There exists a diverse range of deepfake generation techniques, which makes it challenging to develop an accurate universal detection mechanism.
-The 2025 Signal Processing Cup (DFWild-Cup competition) provided a diverse dataset of deepfake images, which are generated from multiple deepfake image generators, for training machine learning model(s) to emphasize the generalization of deepfake detection.
-To this end, we proposed an ensemble-based approach that employs three different neural network architectures: a ResNet-34-based architecture, a data-efficient image transformer (DeiT), and an XceptionNet with Wavelet Transform to capture both local and global features of deepfakes.
-We visualize the specific regions that these models focus for classification using Grad-CAM, and empirically demonstrate the effectiveness of these models in grouping real and fake images into cohesive clusters using t-SNE plots.
-Individually, the ResNet-34 architecture has achieved 88.9% accuracy, whereas the Xception network and the DeiT architecture have achieved 87.76% and 89.32% accuracy, respectively.
-With these networks, our weighted ensemble model achieves an excellent accuracy of 93.23% on the validation dataset of the SP Cup 2025 competition.
-Finally, the confusion matrix and an Area Under the ROC curve of 97.44% further confirm the stability of our proposed method.
+Effective deepfake detection tools are becoming increasingly essential to the growing usage of deepfakes in unethical practices.
+There exists a wide range of deepfake generation techniques, which makes it challenging to develop an accurate universal detection mechanism.
+The 2025 IEEE Signal Processing Cup (DFWild-Cup competition) provided a diverse dataset of deepfake images containing significant class imbalance.
+The images in the dataset are generated from multiple deepfake image generators, for training machine learning model(s) to emphasize the generalization of deepfake detection.
+To this end, we proposed a disjoint set-based multistage training method to address the class imbalance and devised an ensemble-based architecture CAE-Net.
+Our architecture consists of a convolution- and attention-based ensemble network, and employs three different neural network architectures: EfficientNet, Data-Efficient Image Transformer (DeiT), and ConvNeXt with wavelet transform to capture both local and global features of deepfakes.
+We visualize the specific regions that these models focus on for classification using Grad-CAM, and empirically demonstrate the effectiveness of these models in grouping real and fake images into cohesive clusters using t-SNE plots.
+Individually, the EfficientNet B0 architecture has achieved 90.79% accuracy, whereas the ConvNeXt and the DeiT architecture have achieved 89.49% and 89.32% accuracy, respectively.
+With these networks, our weighted ensemble model achieves an excellent accuracy of 94.63% on the validation dataset of the SP Cup 2025 competition.
+The equal error rate of 4.72% and the Area Under the ROC curve of 97.37% further confirm the stability of our proposed method.
+Finally, the robustness of our proposed model against adversarial perturbation attacks is tested as well, showing the inherent defensive properties of the ensemble approach.
 keywords:
-Deepfake , ResNet-34 , Transformer , DeiT , Wavelet , XceptionNet
-journal: Elsevier Image and Vision Computing
+Deepfake detection , ConvNext , Transformer , DeiT , Wavelet , EfficientNet
 […]
@@
 Although deepfake generation tools are widely accessible, detection tools are not nearly as effective and accurate.
-More specifically, the existing detection methods only work well when the test image resembles the features of the images of the training dataset of the detector.
-That is, the major issues in the existing detection approaches are the lack of diversity in training datasets, and the inability to be up to date with new fake image generation techniques [3].
-To address the research gap, the DFWild-Cup competition called for novel approaches that can achieve accurate deepfake detection and provided a diverse dataset (consisting of samples from eight different standard datasets) to help train a generalized deepfake detection algorithm.
-Such a diverse dataset poses unique challenges to formulating accurate image detection methods.
-In our proposed solution for deepfake image detection, we have employed a weighted ensemble approach, which utilized three different neural network architectures with different strengths: i) a ResNet-34-based architecture with squeeze and excitation block, ii) a data-efficient image transformer (DeiT) model, and iii) an Xception network.
-More specifically, we consolidate the outputs from these architectures and employ a weighted ensemble mechanism to get the final result.
-We note that the ResNet-34-based architecture, due to its residual connections, allows learning complex hierarchical features with deeper networks by addressing the problem of vanishing gradients [4].
-Moreover, the squeeze and excitation (SE) block incorporated in the ResNet-34-based architecture provides channel-wise attention with very little computational expense [5].
-Besides, the self-attention mechanism and the distillation token introduced in the DeiT model allow learning long-range dependencies across the entire image.
-More specifically, the ResNet model excels at capturing local spatial relationships of an image, while the transformer model can capture global features of an image through the self-attention mechanism [6].
-Finally, the Xception model, combined with wavelet transform as a preprocessing technique, is effective in detecting fine-grained artifacts in deepfake images.
-We have utilized the ’Haar’ wavelets due to their notable ability to detect abrupt transitions and edges in the data [7].
-A weighted ensemble mechanism of these three networks is used as the final solution, which has enabled us to achieve superior performance compared to the models operating alone.
-Related Works.
-The advancement in deepfake research has led to the creation of nearly perfect image manipulations that are undetectable to the human eye.
+Thus, effective detection of deepfakes is a task of great importance and has been extensively studied for the past decade.
 The combination of multiple deep learning architectures has been shown to perform well on deepfake detection.
-Gowda and Thillaiarasu [8] showed ensemble techniques of Convolutional Neural Network (CNN) models (ResNeXt and Xception) achieve higher accuracy on deepfake detection.
+Gowda and Thillaiarasu [3] showed that ensemble techniques of Convolutional Neural Network (CNN) models (ResNeXt and Xception) achieve higher accuracy on deepfake detection.
 Wodajo and Atnafu [4] combined VGG-16 for feature extraction and Vision Transformer (ViT) for classification, showing promising results on the DFDC dataset.
@@
 uses a spatial-frequency joint dual-stream convolutional neural network with learnable frequency-domain filtering kernels [19].
+Research gap and our contribution.
+The existing detection methods only work well when the test image resembles the features of the images in the training dataset of the detector.
+That is, the major issues in the existing detection approaches are i) the lack of generalizability of extracted features and as such, ii) the inability to be up to date with new fake image generation techniques [20].
+To address the research gap, the DFWild-Cup competition called for novel approaches that can achieve accurate deepfake detection and provided a diverse dataset (consisting of samples from eight different standard datasets) to help train a generalized deepfake detection algorithm.
+Such a diverse dataset poses unique challenges to formulating accurate image detection methods.
+In our proposed solution for generalized deepfake image detection, we have employed a weighted ensemble approach, CAE-Net, which utilizes three different neural network architectures with different strengths: i) an EfficientNet architecture, ii) a Data Efficient Image transformer model, and iii) a ConvNeXt model.
+More specifically, we consolidate the outputs from these architectures and employ a weighted ensemble mechanism to get the final result.
+We note that the EfficientNet-based architecture, due to its depthwise separable convolution and squeeze and excitation block, allows learning fine spatial artifacts and enhances channel-wise attention [21].
+Besides, the self-attention mechanism and the distillation token introduced in the DeiT model allow learning long-range dependencies across the entire image.
+More specifically, the EfficientNet model excels at capturing local spatial relationships of an image, while the transformer model can capture global features of an image through the self-attention mechanism [22].
+Finally, the ConvNeXt model, combined with wavelet transform as a preprocessing technique, is effective in detecting fine-grained artifacts in deepfake images.
+We have utilized the “Haar” wavelets due to their notable ability to detect abrupt transitions and edges in the data [23].
+A weighted ensemble mechanism of these three networks is used as the final solution, which has enabled us to achieve superior performance compared to the models operating alone.
+Finally, we test our models against adversarial perturbation attacks, which shows the trade-offs between weighted ensemble and majority voting ensemble methods.
+The rest of the paper is organized as follows: Section 2 contains data preprocessing techniques, the description of our CAE-Net architecture, consisting of EfficientNet, DeiT, and ConvNeXt.
+Section 3 contains the accuracy results of multiple standard models on the dataset, along with the implementation details and other evaluation metrics of our proposed approach.
+This also contains a comparison of our model’s generalization capabilities against multiple existing works and a defense against adversarial perturbations.
+Finally, Section 4 draws concluding remarks on the contributions presented in the paper.
 2 Proposed Methodology: Weighted Ensemble of Neural Network Models
@@
 The validation dataset consists of 1,548 real images and 1,524 fake images.
-It is evident that there is a significant class imbalance present in the training data as the ratio of real and fake images is approximately 5:1.
-Figure 1: Training procedure including class imbalance handling
+It is evident that there is a significant class imbalance present in the training data, as the ratio of fake and real images is approximately 5:1.
 2.2 Data Preprocessing
 2.2.1 Class Balancing for Model Training
+One of the key contributions of our paper is to devise a disjoint set-based multistage training approach to address the class imbalance present in the given dataset.
 Since the number of fake images is approximately five times the number of real images, we divided the fake images ${}\{\mathcal{F}\}$ into five disjoint subsets $\{\mathcal{F}_{1},\ldots,\mathcal{F}_{5}\}$.
-Each of the three aforementioned models, denoted as $M_{1},M_{2},M_{3}$, was trained in five stages with these disjoint subsets of fake images.
+Each of the three aforementioned models, denoted as $M_{1},M_{2},M_{3}$, are trained in five stages with these disjoint subsets of fake images.
 At each stage $i$, the model is trained using the entire real image set $\mathcal{R}$ and one subset of fake images $\mathcal{F}_{i}$, such that the training dataset of stage $i$ is given by:
@@
 The training of each model is initialized with pre-trained weights $\theta_{0}$, obtained from an ImageNet [32]-trained version of the corresponding architecture.
-The model parameters at each stage are updated iteratively using the dataset $\mathcal{D}_{i}$, following the optimization process:
+The model parameters at each stage are updated iteratively using the dataset $\mathcal{D}_{i}$, through the optimization:
 $\theta_{i}=\arg\min_{\theta}\sum_{(x,y)\in\mathcal{D}_{i}}\mathcal{L}(M(x;\theta),y),$ (2)
@@
 This process continues until the model has been trained on all subsets of fake images, ensuring that the model learns from diverse variations of the fake images while preserving knowledge of real images.
-Figure 1 illustrates the sequential learning procedure.
-2.2.2 Data Augmentation
+2.2.2 Data Augmentation and Normalization
 As mentioned before, we utilized the same set of real images across five training stages to address the class imbalance.
 However, this evidently increases the chance of overfitting.
-To mitigate overfitting and enhance the model’s generalization performance, we implemented four data augmentation techniques: random horizontal flips, random rotations, color jittering, and random resized cropping.
-The random horizontal flip method flips the training images along their vertical axis with a probability of 50%.
-The Random rotation method augments images by randomly rotating them within $0^{o}-10^{o}$.
-The color jitter method randomly alters the brightness, contrast, saturation, and hue of images within a specified range.
-To this end, a range of 0.2 for brightness, contrast, saturation, and hue provided the best results in our experiments.
-Finally, the random resized crop was used as a data augmentation technique that applies both random cropping and resizing to create varied versions of training images.
-In our case, it produced $224\times 224$ images, while randomly scaling the crop area between 80% and 120% of the original image size.
+To mitigate overfitting and enhance the model’s generalization performance, we implemented four data augmentation techniques: random horizontal flips(with 50% probability), random rotations(within $0^{o}-10^{o}$), color jittering (20% variation in brightness, contrast, saturation and hue), and random resized cropping where it produced $224\times 224$ images, while randomly scaling the crop area between 80% and 120% of the original image size.
 These augmentations enhanced the diversity of the training dataset, reduced overfitting, and improved the model’s ability to generalize across a wider range of deepfake detection scenarios.
-It should be noted that such augmentations were used only in training the ResNet-34 and the DeiT architectures.
-For the XceptionNet, we employed wavelet transform-based feature extraction from the raw images, and data augmentation techniques would result in incorrect feature representation in the spectral domain.
-2.2.3 Data Resizing and Normalization
-All three models were pre-trained using the ImageNet dataset.
-Since the requirements of input image dimensions are different on the ResNet-34 architecture and the Xception network, we have resized the input images accordingly to the specific dimensions for respective architectures.
-Additionally, the training images were normalized using the mean and standard deviation of the ImageNet dataset [33].
-2.3 Wavelet Feature Extraction for XceptionNet
+It should be noted that such augmentations were used only in training the DeiT architectures.
+For the ConvNeXt, we employed wavelet transform-based feature extraction from the raw images, and data augmentation techniques would result in incorrect feature representation in the spectral domain.
+Since all three models were pre-trained using the ImageNet dataset, the training images were normalized using the mean and standard deviation of the ImageNet dataset [32].
+The input images were resized accordingly to the specific dimensions for the respective architectures.
+2.3 Wavelet Feature Extraction for ConvNeXt
 The wavelet transform is shown to be effective for image feature extraction as it decomposes images into multi-resolution frequency components [33].
@@
 As such, we performed a single-level 2D Discrete Wavelet Transform (2D-DWT) using Haar wavelets to produce the approximate and detailed coefficients of the images and used the computed coefficients as features.
-Figure 2: Row and Column-wise filtering and downsampling to generate approximation and detail images through Haar 2D-DWT
+Figure 1: Row- and column-wise filtering and downsampling to generate approximation and detail images through 2D-Discrete Wavelet Transform (DWT) using Haar wavelets.
 The 2D-DWT is applied to an image $I(x,y)$ to decompose it into frequency sub-bands while preserving spatial information.
@@
 $F=\text{concat}((A_{r},A_{g},A_{b}),(H_{r},H_{g},H_{b}),(V_{r},V_{g},V_{b}),(D_{r},D_{g},D_{b})).$ (5)
-To prepare the wavelet-domain features for classification, the coefficients are scaled to an 8-bit integer range and resized to match the input dimensions of the XceptionNet model.
+To prepare the wavelet-domain features for classification, the coefficients are scaled to an 8-bit integer range and resized to match the input dimensions of the ConvNeXt model.
 In Figure 2, we showed two such feature images with the approximate and detailed coefficients of two sample images from the dataset.
@@
 The details of the models in our proposed solution are presented below.
-2.4.1 ResNet-34 with SE Block
-We implemented a ResNet-34 model with SE block for the detection of deepfake images.
-Note that, we investigated with different Convolutional Neural Network (CNN)-based architectures – shallow CNNs exhibited high bias, whereas deep CNNs caused vanishing gradients.
-The residual mapping of the ResNet-34 architecture addresses the issue of vanishing gradients, while the capacity of the ResNet-34 architecture is sufficient for capturing the intricate features for deepfake classification.
-Figure 4: Building block of residual learning
-Note that residual learning works by reformulating the target function $\text{H}(x)$ as $\text{F}(x)=\text{H}(x)-x$, where $x$ is the input [4].
-Instead of directly learning $\text{H}(x)$, the network learns the residual $\text{F}(x)$.
-The final output is computed as $\text{F}(x)+x$, combining the learned residual with the input via a skip connection.
-This mechanism allows each layer to focus on learning the residual adjustments needed to improve upon the input.
-The building block of residual mapping is shown in Figure 4.
-Our ResNet-34-based architecture is implemented with two simple principles following the traditional structure of this model: i) When maintaining the same output feature map size, each layer will consistently have the same number of filters.
-ii) Additionally, when the feature map size is halved, the number of filters is effectively doubled.
-Based on these principles, shortcut connections are implemented on a generic CNN architecture to obtain the desired residual network.
-Figure 5: ResNet-34 architecture with Squeeze and Excitation
-Squeeze and Excitation (SE) Block.
-The SE block is introduced as a lightweight attention mechanism to adaptively recalibrate each feature channel’s importance [5].
-The SE block is placed after the final sequential layer (layer 4) of the ResNet-34 model to enhance feature representation by modeling channel-wise dependencies.
+2.4.1 EfficientNet
+EfficientNet-B0 is a computationally efficient model whose architecture is optimized via neural architecture search (NAS) to maximize accuracy under a FLOP constraint, resulting in a highly efficient baseline that outperforms models with similar computational cost [21].
+The core block of the EfficientNet architecture is the Mobile Inverted Bottleneck Convolution (MBConv) block inspired by MobileNetV2 [35].
+The MBConv block consists of a 1x1 Convolution, which expands the input channels to increase feature dimensionality.
+Then, a 3x3 or 5x5 depthwise convolution is applied to capture spatial features efficiently.
+Squeeze-and-Excitation (SE) recalibrates channel importance by computing a channel-wise attention vector [36].
+The SE block acts as a lightweight attention mechanism to adaptively recalibrate each feature channel’s importance [37].
 It is responsible for two main operations:
-  • 1.
 Squeeze: Global spatial information is aggregated using global average pooling to generate channel-wise descriptors.
@@
 where $X_{c}(i,j)$ represents the activation at spatial location $(i,j)$ in channel $c$.
-  • 2.
 Excitation: The descriptors generated from the Squeeze block are passed through a lightweight two-layer fully connected network with a bottleneck structure to model non-linear channel independencies.
-The squeezed feature vector is passed through a gating mechanism with two fully connected layers and a non-linearity (typically ReLU and Sigmoid) to learn channel-wise dependencies:
-    $s=\sigma(W_{2}\delta(W_{1}z)),$ (7)
-    where $W_{1}\in\mathbb{R}^{\frac{C}{r}\times C}$ and $W_{2}\in\mathbb{R}^{C\times\frac{C}{r}}$ are the weights of the two fully connected layers, $\delta(\cdot)$ represents the ReLU activation function, and $\sigma(\cdot)$ is the Sigmoid function that outputs a channel-wise scaling factor $s\in\mathbb{R}^{C}$.
-The final output is obtained by rescaling the input feature maps $X$ using the channel-wise attention weights:
-    $\tilde{X}_{c}=s_{c}X_{c},\quad c=1,2,...,C$ (8)
-    This step selectively enhances informative channels while suppressing less important ones. We denoted this operation as a function $F_{ex}(.)$, as depicted in Fig. 5.
-The details of the ResNet-34 architecture including the output shapes of each layer and the number of parameters are shown in Table 1.
-Table 1: Details of the ResNet-34 Architecture with SE Block
+After that, a skip connection adds the input to the output if the input and output shapes match. The swish activation [38] function is applied after each convolution except the final projection.
+$\text{Swish}(x)=x\cdot\sigma(x)=\frac{x}{1+e^{-x}}$ (7)
+EfficientNet-B0 uses MBConv1 (expansion factor 1) to MBConv6 (expansion factor 6) variants, with kernel sizes of 3x3 or 5x5 depending on the stage.
+A brief summary of different layers, the output shapes and the number of parameters in the EfficientNet-B0 architecture is shown in Table 1.
+Table 1: Summary of the EfficientNetB0 Model Architecture
 Layer (type) Output Shape Param #
-Conv2d, BatchNorm2d, ReLU [64, 112, 112] 9,536
-MaxPool2d [64, 56, 56] 0
-BasicBlock-1 $\times$ 3 [64, 56, 56] 221,952
-   —Conv2d [64, 56, 56] 36,864
-   —BatchNorm2d + ReLU [64, 56, 56] 128
-   —Conv2d [64, 56, 56] 36,864
-   —BatchNorm2d + ReLU [64, 56, 56] 128
-BasicBlock-2 $\times$ 4 [128, 28, 28] 1,116,416
-BasicBlock-3 $\times$ 6 [256, 14, 14] 6,822,400
-BasicBlock-4 $\times$ 3 [512, 7, 7] 13,114,368
-SEBlock (Global Avg Pool + FC) [512, 7, 7] 33,312
-AdaptiveAvgPool2d [512, 1, 1] 0
-Linear [2] 1,026
-Total Parameters 21,319,010
-Trainable Parameters 8,204,642
-Non-Trainable Parameters 13,114,368
-2.4.2 Data Efficient Image Transformer (DeiT) Model
-Unlike conventional Recurrent Neural Networks (RNNs) that process sequential data step by step, transformers process all inputs simultaneously, relying on a self-attention mechanism.
-This simultaneous processing allows transformers to learn long-range dependencies and global contexts of data [6].
-It has been shown that the data-efficient image transformer, which is based on the backbone of the vision transformer, works well for cases where the training data is limited [36].
+Conv2D + BatchNorm + SiLU [32, 112, 112] 928
+DepthwiseSeparableConv [16, 112, 112] 1,448
+InvertedResidual-1 $\times$ 2 [24, 56, 56] 16,714
+InvertedResidual-2 $\times$ 2 [40, 28, 28] 46,640
+InvertedResidual-3 $\times$ 3 [80,14,14] 242.930
+InvertedResidual-4 $\times$ 3 [112, 14, 14] 543,148
+InvertedResidual-5 $\times$ 4 [192, 7, 7] 2,026,348
+InvertedResidual-6 [320, 7, 7] 717,232
+Conv2d + BatchNorm + SiLU [1280, 7, 7] 412,160
+AdaptiveAvgPool2d + Flatten [1280] 0
+Linear [1] 1,281
+Total Parameters 4,008,829
+Trainable Parameters 4,008,829
+Non-Trainable Parameters 0
+2.4.2 Data Efficient Image Transformer (DeiT)
+Unlike conventional Recurrent Neural Networks (RNNs), transformers process all inputs simultaneously, relying on a self-attention mechanism.
+This allows transformers to learn long-range dependencies and global contexts of data [22].
+The data-efficient image transformer, based on the backbone of the vision transformer, has been shown to work well in cases where training data is limited [39].
 Here, an image is divided into a grid of smaller patches as a sequence of input tokens ($N$ number of $16\times 16$ patches), and each patch is associated with a vector through linear embedding.
@@
 Note that the attention mechanism is based on a trainable associative memory of (key, value) pairs.
-A query vector $\mathbf{Q}$ is matched against key vectors $\{\mathbf{K}\}$ using dot products, scaled by $\sqrt{d}$ and normalized via a softmax function to produce weights.
+A query vector $\mathbf{Q}$ is matched against key vectors $\{\mathbf{K}\}$ using dot products, scaled by $\sqrt{d}$, and normalized via a softmax function to produce weights.
 These weights are used to compute a weighted sum of value vectors $\{\mathbf{V}\}$, producing an output matrix as:
-$\text{attention}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\mathbf{V}.$
+$\text{attention}\left(\mathbf{Q},\mathbf{K},\mathbf{V}\right)=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\mathbf{V}$ (8)
 In self-attention, the queries, keys, and values are derived from the same input sequence via linear transformations [41].
@@
 This allows the model to attend to different parts of the input sequence simultaneously.
-Figure 6: Data Efficient Image Transformer (DeiT) architecture
-Note that CNN models primarily capture local contexts and eventually may capture global contexts through hierarchical learning.
-To this end, DeiT is an effective alternative to traditional CNNs in learning global features [6].
-It uses knowledge distillation techniques and optimization strategies to improve generalization.
+DeiT is an effective alternative to traditional CNNs in learning global features [22], using knowledge distillation techniques and optimization strategies to improve generalization.
 DeiT learns from a pre-trained CNN model (that is, the RegNetY-16GF model [42]) and uses a distillation token during training, allowing it to learn faster and more efficiently with less data [39].
@@
 $y_{t}=\text{argmax}_{c}(Z_{t}(c))$ (9)
-Here $c$ is the class index. Using this prediction, the final loss function for hard distillation is treated as
+Here, $c$ is the class index. Using this prediction, the final loss function for hard distillation is treated as
 $L^{\text{hard dist}}_{\text{global}}=\frac{1}{2}L_{CE}(\text{softmax}(Z_{s}),y)+\frac{1}{2}L_{CE}(\text{softmax}(Z_{s}),y_{t})$ (10)
-The first term of the equation can be thought of as $L_{\text{cross-entropy}}$ and the second term as $L_{\text{teacher}}$ depicted in Figure 6
-As the competition dataset lacks enough samples to generalize a ViT, we chose DeiT to capture global contexts effectively, despite it being computationally costly.
-There exist different DeiT model variations based on the number of heads and embedding resolutions.
-In our proposed solution, we have implemented the DeiT-B model, which uses the ViT-B model as its backbone.
-It has 768 embedding dimensions with 12 heads containing a total of 86 million parameters [36].
-The DeiT model architecture is shown in Figure 6 and the number of parameters and the output shapes of each layer are shown in Table 2.
-Table 2: Details of the Data Efficient Image Transformer (DeiT) Model
+The first term of the equation can be thought of as $L_{\text{cross-entropy}}$ and the second term as $L_{\text{teacher}}$.
+Our proposed solution utilizes the DeiT-B model, a 768-dimensional model with 12 heads and 86 million parameters, based on the ViT-B model [39].
+The summary number of parameters and the output shapes of layers are shown in Table 2.
+Table 2: Summary of the Data Efficient Image Transformer (DeiT) Model Architecture
 Layer (type) Output Shape Param #
@@
 Patch Embed [196, 768] 0
-Dropout [197, 768] 0
-Identity-1 [197, 768] 0
+Dropout + Identity-1 [197, 768] 0
 LayerNorm $\times$ 25 [197, 768] 38,400
@@
 Identity-2 [12, 197, 64] 0
-Linear-2 $\times$ 12 [197, 768] 35,407,872
-Attention $\times$ 12 [197, 768] 0
-Linear-3 $\times$ 12 [197, 3072] 28,348,416
-GELU $\times$ 12 [197, 3072] 0
+(Linear-2 + Attention) $\times$ 12 [197, 768] 35,407,872
+(Linear-3 + GELU) $\times$ 12 [197, 3072] 28,348,416
 Dropout [197, 3072] 0
@@
 Non-Trainable Parameters 0
-2.4.3 XceptionNet Model
-XceptionNet is based on the concept of depthwise separable convolutions, which enhance computational efficiency while maintaining accuracy.
-In depthwise convolution, a single convolutional filter is applied to each input channel.
-Following this, pointwise convolution is used to combine the outputs of the depthwise convolution into a linear form [40].
-XceptionNet features a fully convolutional architecture for simplicity and efficiency, leverages residual connections to improve gradient flow, and employs deep feature extraction.
-(a) Entry Flow
-(b) Middle Flow
-(c) Exit Flow
-Figure 7: Details of XceptionNet architecture
-As shown in Figure 7, our implemented XceptionNet has 36 convolutional layers, structured into 14 modules with residual connections, except for the first and last modules [40].
-It has three primary sections—Entry Flow, Middle Flow, and Exit Flow.
-The Entry Flow layer consists of convolutional layers followed by max pooling and is responsible for reducing spatial dimensions and extracting low-level features.
-The Middle Flow layer contains multiple depthwise convolution blocks with residual connection, focusing on high-level features.
-The Exit Flow layer includes more depthwise separable convolutions and reduces spatial dimensions further.
-We modified the fully connected layer from the original architecture to match our binary classification problem.
-The summary of the output shapes and number of parameters of each block and layer is shown in Table 3.
-Table 3: Details of the XceptionNet Model
-Layer (type) Output Shape Param #
-Conv2d [32, 149, 149] 864
-BatchNorm2d + ReLU [32, 149, 149] 64
-Conv2d [64, 147, 147] 18,432
-BatchNorm2d + ReLU [64, 147, 147] 128
-Block-1 [128, 74, 74] 35,264
-   —SeparableConv2d [128, 147, 147] 8,768
-   —BatchNorm2d + ReLU [128, 147, 147] 256
-   —SeparableConv2d [128, 147, 147] 17,536
-   —BatchNorm2d [128, 147, 147] 256
-   —MaxPool2d [128, 74, 74] 0
-   —Conv2d [128, 74, 74] 8,192
-   —BatchNorm2d [128, 74, 74] 256
-Block-2 [256, 37, 37] 136,064
-Block-3 [728, 19, 19] 915,944
-Block-4 $\times$ 8 [728, 19, 19] 12,911,808
-Block-5 [1024, 10, 10] 2,039.584
-SeparableConv2d [1536, 10, 10] 1,582,080
-BatchNorm2d + ReLU [1536, 10, 10] 3072
-SeparableConv2d [2048, 10, 10] 3,159,552
-BatchNorm2d + ReLU [2048, 10, 10] 4096
-SelectAdaptivePool2d [2048] 0
-Linear [2] 4098
-Total Parameters 20,811,050
-Trainable Parameters 20,811,050
+2.4.3 ConvNeXt Model
+ConvNeXt combines the efficiency and inductive biases of traditional CNNs with the scalability and performance of transformers.
+It employs an inverted bottleneck block, similar to transformers’ MLP blocks, reducing FLOPs in downsampling layers.
+Here, batch normalization is replaced with layer normalization, aligning with transformer practices, and GELU is used instead of ReLU for smoother non-linearity [43].
+Inspired by both ResNets [44] and Swin transformers [45], the architecture includes a patchify stem where input images are processed by a 4x4 convolution with stride 4, followed by layer normalization, to downsample the image into non-overlapping patches.
+This mimics ViT’s patch embedding but uses convolutions.
+Each ConvNeXt block contains a 7x7 depthwise convolution (one filter per input channel) that is applied to capture spatial features efficiently.
+Then layer normalization is applied before and after the convolution to stabilize training.
+It normalizes features across the channel dimension for each spatial position.
+For feature $X[i,j,:]~{}\epsilon~{}\mathbf{R^{C}}$ the normalized output is:
+$\hat{X}[i,j,c]=\frac{X[i,j,c]-\mu}{\sqrt{\sigma^{2}+\epsilon}}\cdot\gamma+\beta$ (11)
+Here $\mu=\frac{1}{C}\sum_{c=1}^{C}X[i,j,c]$, $\sigma^{2}=\frac{1}{C}\sum_{c=1}^{C}(X[i,j,c]-\mu)^{2}$ , $\epsilon$ is a small constant, and $\gamma$ and $\beta$ are trainable parameters.
+This type of normalization works better than typical batch normalization [46].
+After that, pointwise convolution is applied to expand (to 4x the input dimension) and contract the feature map, mimicking transformer MLP layers.
+GELU (Gaussian Error Linear Unit) [47] is applied instead of RELU for smoother non-linearity.
+$\text{GELU}(x)=x\cdot\Phi(x)=x\cdot\frac{1}{2}[1+\text{erf}(\frac{x}{\sqrt{2}})]$ (12)
+A skip connection is used from the input to the block’s output, similar to ResNet’s residual connections [44].
+Between stages, ConvNeXt uses a downsampling layer with layer normalization and a 2x2 convolution with stride 2, inspired by Swin Transformer’s separate downsampling strategy, unlike ResNet’s within-block downsampling.
+From the diverse family of ConvNeXt architectures, we specifically implemented the ConvNeXt-tiny model for its low compute and high accuracy on the dataset.
+A summary of the different layers in the ConvNeXt-tiny architecture that we have implemented is presented in Table 3.
+Table 3: Summary of the ConvNeXt Architecture
+Layer Type Output Shape Param #
+Conv2d + LayerNorm2d $[96,74,74]$ 4,896
+CNBlock (×$\times 3$) $[96,74,74]$ 237,600
+LayerNorm2d + Conv2d $[192,37,37]$ 74,112
+CNBlock (×$\times 3$) $[192,37,37]$ 917,568
+LayerNorm2d + Conv2d $[384,18,18]$ 295,680
+CNBlock (×$\times 9$) $[384,18,18]$ 10,813,824
+LayerNorm2d + Conv2d $[768,9,9]$ 1,181,184
+CNBlock (×$\times 3$) $[768,9,9]$ 14,287,104
+AdaptiveAvgPool2d + LayerNorm2d $[768,1,1]$ 1536
+Flatten + Linear $[1]$ 769
+Total Parameters 27,814,273
+Trainable Parameters 27,814,273
 Non-Trainable Parameters 0
-2.4.4 Weighted Ensemble Mechanism
+2.4.4 Proposed Weighted Ensemble Architecture: CAE-Net
 Finally, we implemented a weighted ensemble mechanism to combine the predictions of the aforementioned three models.
-The associated weights for the ResNet-34 with SE, DeiT, and Xception with wavelet transform are 0.36, 0.45, and 0.19, respectively, based on our parameter search.
-The final architecture is shown in Figure 8, and the total number of trainable and non-trainable parameters in the final architecture is shown in Table 4.
-Figure 8: The proposed weighted ensemble mechanism
-Table 4: Number of parameters in the final weighted ensemble mechanism
-Model Total Parameters Trainable Non-trainable
-ResNet 21,319,010 8,204,642 13,114,368
-DeiT 85,648,130 85,648,130 0
-Xception 20,811,050 20,811,050 0
-Total 127,778,990 114,664,622 13,114,368
+The associated weights for the EfficientNet, DeiT, and ConvNeXt with wavelet transform are 0.35, 0.34, and 0.31, respectively, based on our parameter search.
+The final architecture is shown in Figure 3.
+Figure 3: The proposed CAE-Net architecture
+2.5 Adversarial Perturbations
+Adversarial perturbations are crafted noises or modifications added to classifier model inputs, often imperceptible to humans, that can significantly deteriorate model performance and lead to misclassification.
+Such adversarial perturbations can be universal or image-dependent [48] and are usually done in white-box settings where access to the model’s architecture and parameters is required.
+Also, adversarial attacks on a known network have been shown to work effectively against other networks as well [49], making black-box attacks an imminent threat for sensitive issues like deepfake detection [50].
+We evaluate the robustness of our proposed models by testing them against a simple but effective adversarial perturbation, the Fast Sign Gradient Method (FGSM).
+FGSM is a white-box attack without specifying a target class.
+Using this method, adversarial images are created by calculating the gradient of loss for input data and adjusting the input by a small amount in the direction of the gradient sign [49].
+$x_{\text{adv}}=x+\epsilon\cdot\text{sign}(\nabla_{x}J(\theta,x,y))$ (13)
+Here $(x,y)$ are the original input image and its label, $\theta$ indicates model parameters, $x_{adv}$ is the adversarially perturbed image, and $\nabla_{x}J(\theta,x,y)$ is the gradient of the loss function $J$ with respect to the input image $x$.
+The parameter $\epsilon$ controls the magnitude of the perturbation added to the original image.
+This method works like a gradient ascent algorithm trying to maximize the loss with the smallest possible perturbation, $\epsilon$, which makes it barely noticeable to the human eye.
+Though FGSM is one of the simplest forms of perturbations, it has been shown to reduce the accuracy of high-quality detectors to a large extent [51].
+Its effectiveness for binary classification problems makes it an ideal choice for testing the robustness of our proposed method.
 3 Experimental Results
-Figure 9: (a) Confusion matrix and (b) ROC curve for validation set
 In this section, we demonstrate the performance of our proposed solution on the competition dataset.
-We note that investigation and experimentation with different architectures helped us make informed decisions to reach our proposed solution.
-More specifically, we started experimenting with simpler CNN models.
-As mentioned before, the performance of CNN models, such as the VGG-19, indicated insufficient capacity and indicated the necessity of deeper networks.
-Additionally, residual networks, such as the ResNet-50, raised overfitting issues.
-These results motivated us to choose a ResNet-34-based architecture to strike a balance between high-bias and high-variance regimes.
-We investigated the performance of the convolutional block attention module (CBAM) and SE and opted for SE block based on the generalization performance on the given dataset.
-Lastly, we investigated the effect of different handcrafted feature extraction methods on the performance and opted for wavelet feature extraction for our XceptionNet model.
-(a) SEResNet (Untrained)
-(b) DeiT (Untrained)
-(c) Xception (Untrained)
-(d) SEResNet (Trained)
-(e) DeiT (Trained)
-(f) Xception (Trained)
-Figure 10: t-SNE plot of three models before and after training for randomly selected 4000 real and fake images using perplexity = 40
-In essence, gradually increasing the complexity of networks helped us understand the shortcomings of different architectures for our specified task and led us to choose the model architectures based on their generalization capability with a relatively small dataset.
-To further reduce the variance, we employed a weighted ensemble of ResNet-34 with SE, DeiT, and XceptionNet to generate the final results.
-We present a comparative analysis of the results of different model variations and configurations that guided us to our final design decisions in Table.
-5.
-We argue that these results serve as empirical evidence of our choice of neural network architectures, and the weighted ensemble approach.
+No prior work exists on this novel combined deepfake image detection dataset, which consists of images from eight different benchmark deepfake video datasets.
+To address this, we hypothesized that lightweight CNNs(e.g., EfficientNet) would excel in resource-constrained settings, modern CNNs (e.g., ConvNeXt) would leverage transformer-inspired designs for better feature extraction, and ViTs would capture the global context of images.
+We systematically evaluated these models to test these hypotheses along with multiple other frequency domain feature extraction techniques, establishing the first performance benchmarks for the dataset.
 Table 4: Comparative Analysis of Different Model Variations
-Model Validation Acc.
-Xception-Net [40] 61.48%
+Classifier Models Accuracy
+CNN models
   VGG-19 [52] 62.25%
   ResNet-50 [44] 69.00%
+  MobileNet-V2 78.12%
+  ResNet-34 [44] 80.32%
+  RegNet [53] 80.92%
+  Inception V3 [54] 84.38%
+  ResNet-34 with CBAM 85.19%
+  ResNet-34 with single SE 88.90%
+  DenseNet-121 [55] 90.01%
+  EfficientNet-B0 [21] 90.79%
+  EfficientNet-B4 [21] 87.27%
+  EfficientNet-B5 [21] 88.31%
+  ResNeXt [56] 85.19%
+  ConvNeXt [43] 88.18%
+Frequency Domain Experimentations
   DCT with ResNet-34 71.48%
-ResNet-34 [4] 80.32%
-ResNet-34 with CBAM 85.19%
-Xception-Net with Wavelet Transform 87.76%
-ResNet-34 with SE 88.90%
+  ResNet-34 with Wavelet 80.44%
+  EfficientNetB0 with Wavelet 82.48%
+  Xception-Net [57] with Wavelet 87.76%
+  ConvNeXt with Wavelet 89.49%
+Study on Transformers
+  ViT-B [58] 78.12%
+  Swin Transformer [45] 86.52%
+  MobileViT [59] 87.04%
   DeiT [39] 89.32%
-Maj. Voting (ResNet, DeiT, Xception) 91.34%
-Weighted Ensem. (ResNet, DeiT) 91.80%
-Weighted Ensem. (ResNet, DeiT & Xception) (Proposed) 93.23%
+Ensemble Results
+  Maj. Voting (EfficientNet, DeiT, ConvNeXt) 94.24%
+  Proposed Weighted Ensem. (EfficientNet, DeiT, ConvNeXt) 94.63%
 Hyperparameters.
-For the training of our models, we have used the ADAM optimizer.
-We have used accuracy, precision, recall, and F1 score as our evaluation metrics.
-Since the deepfake detection problem is a binary classification problem, we used binary categorical cross-entropy as our loss function.
-Using the validation set to observe the accuracy of the proposed architecture, each image took approximately 0.03515625 sec on a Tesla P100 GPU.
+For the training of our models, we have used the ADAM optimizer and binary categorical cross-entropy as our loss function.
+Using the validation set to observe the accuracy of the proposed architecture, each image took approximately 0.03385416s on a Tesla P100 GPU.
 The hyperparameters used for training are presented in Table 5.
@@
 Batch size 32
-Number of Epochs
-   - ResNet-34 with SE 10
+Number of Epochs in each session
+   - EfficientNet 5
    - DeiT 5
-   - XceptionNet 10
+   - ConvNeXt 4
 Learning rate
@@
    - Patience 3
-We have utilized the NumPy and Pandas libraries for preprocessing and data manipulation tasks, and PyTorch [42] for model implementation, training, and validation.
-More specifically, we used torch for core functionality, torch.nn for neural networks, torch.optim for optimization, and utilities like torch.utils.data.Dataset, DataLoader, and random_split.
-From TorchVision, we used transforms for image transformations and models for pre-trained architectures.
-For performance metrics, we applied Scikit-learn (precision_score, recall_score, and f1_score).
-Progress tracking was managed with TQDM, and TorchSummary provided model summaries.
-Image processing utilities include OpenCV, while PyWavelets was utilized for wavelet transformations.
-Additionally, Timm was used to access advanced pre-trained models, and the Random module was applied for generating randomness.
-(a) SEResNet
-(b) DeiT
-(c) Xception
-(d) SEResNet
-(e) DeiT
-(f) Xception
-Figure 11: Importance map shown by Grad-CAM for different models
-(a) SEResNet
-(b) DeiT
-(c) Xception
-Figure 12: Grad-CAM view of a few failed classifications of our proposed model
 Result Analysis.
-As mentioned before, we evaluated our weighted ensemble model for multiple evaluation metrics, including accuracy, precision, recall, F1-score, area under receiver operating characteristic (AU-ROC) curve, equal error rate, and average precision.
-While calculating these metrics, fake images were considered as class 0 and real images as class 1.
+We evaluated our weighted ensemble model for multiple evaluation metrics, including accuracy, precision, recall, F1-score, area under the receiver operating characteristic (AU-ROC) curve, equal error rate, and average precision.
+While calculating these metrics, real images were considered as class 0 and fake images as class 1.
 The performance of our model on these metrics is presented in Table 6 and the confusion matrix and the ROC curve are shown in Figure 4.
 It is evident from the evaluation indices that our proposed weighted ensemble architecture performs very well on the competition dataset, which underlines the generalization capability of the proposed approach.
-Table 7: Performance of the Final Ensemble Architecture combining ResNet-34, DeiT, and XceptionNet
+Table 6: Performance of the Final Ensemble Architecture combining ConvNext, DeiT, and EfficientNet
 Evaluation Metrics Performance
-Accuracy 93.23%
-Precision 0.9515
-Recall 0.9121
-F1 score 0.9314
-Area under ROC 0.9769
-Equal Error Rate 6.33%
-Average Precision 0.9744
-To investigate the performance of our ensemble model, we intend to visualize i) the regions of the images that the three models focus for classification, and ii) how the three models distinguish real and fake images before and after training.
+Accuracy 94.63%
+Precision 0.937
+Recall 0.956
+F1 score 0.9464
+Area under ROC 0.9737
+Equal Error Rate 4.72%
+Average Precision 0.9771
+Figure 4: (a) Confusion matrix and (b) ROC curve for validation set
+To investigate the performance of our ensemble model, we intend to visualize i) the regions of the images that the three models focus on for classification, and ii) how the three models distinguish real and fake images before and after training.
 To that end, we note that the Grad-CAM visualizations show how different models focus on different facial regions of the same image for deepfake detection.
-In Figure 11, it can be seen that the ResNet-34 and Xception models focus on localized facial features, whereas DeiT displays diffuse attention patterns indicating a focus on global features.
-Each of the models focuses on different parts of the image, capturing unique and complementary information.
-When these diverse insights are combined into an ensemble, the overall prediction becomes more accurate, robust, and better at generalizing to new data.
-Additionally, we present the Grad-CAM view of some of the wrong predictions by the three models in Figure 12, showing the focus of irrelevant regions, leading to mis-classifications.
+In Figure 5, it can be seen that EfficientNet focuses on localized facial features, whereas DeiT displays diffuse attention patterns indicating a focus on global features.
+When the diverse insights from each model are combined into an ensemble, the overall prediction becomes more accurate, robust, and better at generalizing to new data.
+Additionally, we present the Grad-CAM view of some of the wrong predictions by the three models in Figure 6, showing the focus of irrelevant regions, leading to misclassifications.
+(a) EfficientNet
+(b) DeiT
+(c) ConvNeXt
+Figure 5: Importance map shown by Grad-CAM for different models
+(a) EfficientNet
+(b) DeiT
+(c) ConvNeXt
+Figure 6: Grad-CAM view of a few failed classifications of our proposed model
+(a) EfficientNet (Untrained)
+(b) DeiT (Untrained)
+(c) ConvNeXt (Untrained)
+(d) EfficientNet (Trained)
+(e) DeiT (Trained)
+(f) ConvNeXt (Trained)
+Figure 7: t-SNE plot of three models before and after training for randomly selected 4000 real and fake images using perplexity = 40
 Next, we note that high-dimensional data can be visualized by projecting onto a two-dimensional space using the non-linear dimensionality reduction technique known as t-SNE (t-Distributed Stochastic Neighbor Embedding) [60].
@@
 As such, the t-SNE visualizations (Fig.
-10) offer valuable insights into the effectiveness of our models.
-For both real and fake images, we observe a single cohesive cluster for all three models.
-Notably, the cluster of fake images generated by two of our models, SEResNet and XceptionNet, is compact, despite these images being produced using various deepfake generation methods.
-The plots show that the training has effectively enhanced the ability to distinguish between real and fake image features for all three models.
-This, in-turn, highlights the generalization capability of our proposed ensemble model.
-After training, all three backbone models—SEResNet, DeiT, and XceptionNet—produced well-separated embeddings for real and fake images, as illustrated in Fig.
-10(d), 10(e), and 10(f).
+7) offer valuable insights into the effectiveness of our models.
+For both real and fake images, we observe cohesive clusters for all three models, despite the images being produced using various deepfake generation methods.
+This highlights the generalization capability of our proposed ensemble model.
+After training, all three backbone models, i.e., EfficientNet, DeiT, and ConvNeXt, produced well-separated embeddings for real and fake images, as illustrated in Fig.
+7(d), 7(e), and 7(f).
+There exist very few works on deepfake detection that perform well on the combination of such a diverse set of datasets.
+A comparative analysis among existing works with our proposed architecture is shown in Table 7.
+Note that the AUC score presented here is the mean score of these models on the corresponding datasets (we refer the reader to Figure 8 of [25]).
+We want to emphasize that existing works often demonstrate good performance on only a few datasets, but the overall performance on multiple datasets is very poor for most architectures.
+Table 7: Comparative Analysis of Average AUC Score of Different Models
+[…]
+[…]
+DFD [27], DFDC [31], Celeb-DF [25]
+Meso4 [63] UADFV [29], DF-TIMIT [62], FaceForensics++ [26], 75.9% 10.63%
+DFD [27], DFDC [31], Celeb-DF [25]
+FWA [64] UADFV [29], DF-TIMIT [62], FaceForensics++ [26], 82.1% 14.45%
+DFD [27], DFDC [31], Celeb-DF [25]
+Xception-C23 [26] UADFV [29], DF-TIMIT [62], FaceForensics++ [26], 86.4% 11.96%
+DFD [27], DFDC [31], Celeb-DF [25]
+DSP-FWA [25] UADFV [29], DF-TIMIT [62], FaceForensics++ [26], 87.4% 12.79%
+DFD [27], DFDC [31], Celeb-DF [25]
+Our Proposed Method Celeb-DF [25], DFDC [31], DFD [27], UADFV [29], 97.37% -
+FaceShifter [28], DFDC Preview [30], FaceForensics++ [26]
+Finally, we look into the effectiveness of our ensemble method through a simple adversarial perturbation attack, such as Fast Gradien Sign Method (FGSM).
+As mentioned before, FGSM can fool models by adding simple perturbation noise based on the gradient sign that intends to maximize the loss function of the detection model instead of minimizing it.
+We present the accuracy scores of the separate detection model as well as the ensemble models for different values of $\epsilon$ in Table 8.
+Table 8: Comparison of Our Proposed Architectures under Adversarial FGSM Attacks
+Model Accuracy
+$\epsilon=0.01$ $\epsilon=0.03$ $\epsilon=0.05$
+EfficientNetB0 38.54% 46.48% 53.94%
+Data Efficient Image Transformer 42.76% 21.94% 16.73%
+ConvNeXt with wavelet 6.45% 6.51% 15.40%
+Weighted Ensem.(EfficientNet, DeiT, ConvNeXt) 12.04% 13.48% 16.05%
+Majority Voting (EfficientNet, DeiT, ConvNeXt) 25.52% 20.41% 19.76%
+The results show an increased robustness of the EfficientNet architecture, which depicts increased accuracy with increasing perturbation compared to the other two architectures.
+DeiT shows a decrease in accuracy with rising perturbation, and ConvNeXt accuracy drops significantly even for very low perturbation added to the image.
+Though a weighted ensemble model is appropriate for achieving high accuracy on the generalized dataset, we argue that it is highly vulnerable to adversarial attacks.
+If any model performs poorly on adversarial data, then the performance of the entire ensemble drops significantly.
+However, in the majority voting ensemble, the performance does not reduce as much, since it requires the majority of models to perform poorly, implying increased robustness against adversarial attacks for applications, such as deepfake detection.
 4 Conclusion
-In this work, we have proposed a weighted ensemble approach that takes advantage of the strengths of a ResNet-34-based model, a DeiT model, and an XceptionNet model for the accurate classification of deepfake images.
-The ResNet-34 model enabled extracting local features from an image, while the DeiT model specialized in capturing global features.
-Additionally, XceptionNet, combined with wavelet transform, effectively identified frequency-dependent artifacts.
+In this work, we proposed an ensemble architecture for deepfake detection that captured the intricate properties of the input images to successfully distinguish between real and fake images and produced excellent results.
 We emphasize that detecting deepfake images that are generated with a range of generative AI approaches is a challenging task.
-Our proposed model architectures captured the intricate properties of the input images to successfully distinguish between real and fake images and produced excellent results.
+To that end, we have addressed a significant class imbalance through disjoint training set-based multistage training, and proposed a weighted ensemble approach that takes advantage of the strengths of an EfficientNet model, a Data Efficient Image Transformer model, and a ConvNext model for the accurate classification of deepfake images.
+The EfficientNet-B0 model enabled extracting local features from an image, while the DeiT model specialized in capturing global features.
+Additionally, ConvNeXt, combined with the wavelet transform, effectively identified frequency-dependent artifacts.
 Through t-SNE plots, we demonstrated that our model is capable of separating real and fake images into separable clusters.
-Additionally, we included a Grad-CAM visualizations to map the focus-regions of images to illustrate how our models analyze the images and which portions contribute to distinguishing between real and fake images.
-An interesting future work could be the incorporation of lightweight architectures and reduce the computational overhead, as well as an increase in the model’s adaptability to newer and sophisticated deepfake generation methods.
+Additionally, we included Grad-CAM visualizations to map the focus regions of images to illustrate how our models analyze the images and which portions contribute to distinguishing between real and fake images.
+Besides, we further analyzed the effectiveness of the proposed models by testing them against FGSM attacks and found the optimum ensemble mechanism to be majority voting compared to the weighted ensemble with a little trade-off in accuracy.
+An interesting future work could be incorporating lightweight architectures, reducing the computational overhead, and increasing the model’s adaptability to newer and sophisticated deepfake generation methods.

### v3

@@
 Abstract
-Effective deepfake detection tools are becoming increasingly essential to the growing usage of deepfakes in unethical practices.
-There exists a wide range of deepfake generation techniques, which makes it challenging to develop an accurate universal detection mechanism.
-The 2025 IEEE Signal Processing Cup (DFWild-Cup competition) provided a diverse dataset of deepfake images containing significant class imbalance.
-The images in the dataset are generated from multiple deepfake image generators, for training machine learning model(s) to emphasize the generalization of deepfake detection.
-To this end, we proposed a disjoint set-based multistage training method to address the class imbalance and devised an ensemble-based architecture CAE-Net.
-Our architecture consists of a convolution- and attention-based ensemble network, and employs three different neural network architectures: EfficientNet, Data-Efficient Image Transformer (DeiT), and ConvNeXt with wavelet transform to capture both local and global features of deepfakes.
-We visualize the specific regions that these models focus on for classification using Grad-CAM, and empirically demonstrate the effectiveness of these models in grouping real and fake images into cohesive clusters using t-SNE plots.
-Individually, the EfficientNet B0 architecture has achieved 90.79% accuracy, whereas the ConvNeXt and the DeiT architecture have achieved 89.49% and 89.32% accuracy, respectively.
-With these networks, our weighted ensemble model achieves an excellent accuracy of 94.63% on the validation dataset of the SP Cup 2025 competition.
-The equal error rate of 4.72% and the Area Under the ROC curve of 97.37% further confirm the stability of our proposed method.
-Finally, the robustness of our proposed model against adversarial perturbation attacks is tested as well, showing the inherent defensive properties of the ensemble approach.
-keywords:
-Deepfake detection , ConvNext , Transformer , DeiT , Wavelet , EfficientNet
-[…]
-[inst1]organization=Department of Electrical and Electronic Engineering,addressline=BUET, city=Dhaka, postcode=1205, country=Bangladesh
-[…]
-[inst2]organization=School of Data and Sciences,addressline=BRAC University, city=Dhaka, postcode=1212, country=Bangladesh
+The spread of deepfakes poses significant security concerns, demanding reliable detection methods.
+However, diverse generation techniques and class imbalance in datasets create challenges.
+We propose CAE-Net, a Convolution- and Attention-based weighted Ensemble network combining spatial and frequency-domain features for effective deepfake detection.
+The architecture integrates EfficientNet, Data-Efficient Image Transformer (DeiT), and ConvNeXt with wavelet features to learn complementary representations.
+We evaluated CAE-Net on the diverse IEEE Signal Processing Cup 2025 (DF-Wild Cup) dataset, which has a 5:1 fake-to-real class imbalance.
+To address this, we introduce a multistage disjoint-subset training strategy, sequentially training the model on non-overlapping subsets of the fake class while retaining knowledge across stages.
+Our approach achieved $94.46\%$ accuracy and a $97.60\%$ AUC, outperforming conventional class-balancing methods.
+Visualizations confirm the network focuses on meaningful facial regions, and our ensemble design demonstrates robustness against adversarial attacks, positioning CAE-Net as a dependable and generalized deepfake detection framework.
+This manuscript has been published in the Journal of Visual Communication and Image Representation by Academic Press.
+Keywords: Deepfake detection, ConvNext, Transformer, DeiT, Wavelet, EfficientNet
 1 Introduction
@@
 The combination of multiple deep learning architectures has been shown to perform well on deepfake detection.
-Gowda and Thillaiarasu [3] showed that ensemble techniques of Convolutional Neural Network (CNN) models (ResNeXt and Xception) achieve higher accuracy on deepfake detection.
-Wodajo and Atnafu [4] combined VGG-16 for feature extraction and Vision Transformer (ViT) for classification, showing promising results on the DFDC dataset.
+Gowda and Thillaiarasu [3] showed that ensemble techniques of Convolutional Neural Network (CNN) models (ResNeXt [4] and Xception [5]) achieve higher accuracy on deepfake detection.
+Wodajo and Atnafu [6] combined VGG-16 [7] for feature extraction and Vision Transformer (ViT) [8] for classification, showing promising results on the DFDC dataset.
 Wolter et al.
-[5] proposed a wavelet-packet-based analysis for detecting GAN-generated images, highlighting spatial frequency differences in synthetic content and achieving competitive results using combined architectures of wavelets and CNNs.
+[9] proposed a wavelet-packet-based analysis for detecting Generative Adversarial Networks (GAN)-generated images, highlighting spatial frequency differences in synthetic content and achieving competitive results using combined architectures of wavelets and CNNs.
 Ricker et al.
-[6] proposed the usage of frequency domain features extracted via DCT, and achieved high accuracy rates (97.7% for GAN and 73% for diffusion models), highlighting the utility of frequency-domain artifacts.
+[10] proposed the usage of frequency domain features extracted via Discrete Cosine Transform (DCT) [11], and achieved high accuracy rates (97.7% for GAN and 73% for diffusion models), highlighting the utility of frequency-domain artifacts.
 Patel et al.
@@
 However, Abdullah et al.
-[9] highlighted two major limitations of the existing networks arguing that the state-of-the-art models often lack control over content and image quality, making them ineffective at detecting high-quality deepfakes if trained on low-quality fake images, and they also lack adversarial evaluation, where an attacker can exploit knowledge of the defense.
+[14] highlighted two major limitations of the existing networks, arguing that the state-of-the-art models often lack control over content and image quality, making them ineffective at detecting high-quality deepfakes if trained on low-quality fake images, and they also lack adversarial evaluation, where an attacker can exploit knowledge of the defense.
 Works on adversarial perturbation attacks to fool deep neural network (DNN) based detectors are a challenge in deepfake detection [15, 16].
@@
 Lin et al.
-proposed a CNN-based method for deepfake detection that combines multi-scale convolution with a vision transformer [13].
+proposed a CNN-based method for deepfake detection that combines multi-scale convolution with a vision transformer (ViT) [18].
 Their approach features a multi-scale module incorporating dilation convolution and depth-wise separable convolution.
@@
 Research gap and our contribution.
-The existing detection methods only work well when the test image resembles the features of the images in the training dataset of the detector.
-That is, the major issues in the existing detection approaches are i) the lack of generalizability of extracted features and as such, ii) the inability to be up to date with new fake image generation techniques [20].
-To address the research gap, the DFWild-Cup competition called for novel approaches that can achieve accurate deepfake detection and provided a diverse dataset (consisting of samples from eight different standard datasets) to help train a generalized deepfake detection algorithm.
-Such a diverse dataset poses unique challenges to formulating accurate image detection methods.
-In our proposed solution for generalized deepfake image detection, we have employed a weighted ensemble approach, CAE-Net, which utilizes three different neural network architectures with different strengths: i) an EfficientNet architecture, ii) a Data Efficient Image transformer model, and iii) a ConvNeXt model.
-More specifically, we consolidate the outputs from these architectures and employ a weighted ensemble mechanism to get the final result.
-We note that the EfficientNet-based architecture, due to its depthwise separable convolution and squeeze and excitation block, allows learning fine spatial artifacts and enhances channel-wise attention [21].
-Besides, the self-attention mechanism and the distillation token introduced in the DeiT model allow learning long-range dependencies across the entire image.
-More specifically, the EfficientNet model excels at capturing local spatial relationships of an image, while the transformer model can capture global features of an image through the self-attention mechanism [22].
-Finally, the ConvNeXt model, combined with wavelet transform as a preprocessing technique, is effective in detecting fine-grained artifacts in deepfake images.
-We have utilized the “Haar” wavelets due to their notable ability to detect abrupt transitions and edges in the data [23].
-A weighted ensemble mechanism of these three networks is used as the final solution, which has enabled us to achieve superior performance compared to the models operating alone.
-Finally, we test our models against adversarial perturbation attacks, which shows the trade-offs between weighted ensemble and majority voting ensemble methods.
-The rest of the paper is organized as follows: Section 2 contains data preprocessing techniques, the description of our CAE-Net architecture, consisting of EfficientNet, DeiT, and ConvNeXt.
-Section 3 contains the accuracy results of multiple standard models on the dataset, along with the implementation details and other evaluation metrics of our proposed approach.
+We note that the existing detection methods only work well when the test image resembles the features of the images in the training dataset of the detector.
+As such, the major issues in the existing detection approaches are i) the lack of generalizability of extracted features, and ii) the inability to be up to date with new fake image generation techniques [25].
+Motivated by this limitation, we propose a weighted ensemble-based deepfake detection framework, CAE-Net, which combines convolutional and attention mechanisms to capture both spatial and frequency-domain cues.
+The model integrates three different neural network architectures with complementary strengths: i) an EfficientNet architecture [26], ii) a Data-Efficient Image Transformer (DeiT) [27], and iii) a ConvNeXt model [28] enhanced with a wavelet-based preprocessing stage.
+This hybrid configuration, to the best of our knowledge, is the first to incorporate wavelet-domain feature enhancement within a CNN–Transformer ensemble, effectively leveraging local, global, and frequency-domain representations for generalized deepfake image detection.
+Since training on highly diverse datasets remains difficult for most existing architectures, we employ the IEEE Signal Processing Cup 2025 (DF-Wild Cup) dataset to demonstrate our proposed CAE-Net’s generalization capabilities.
+This dataset combines eight publicly available deepfake benchmark datasets under the DeepfakeBench framework and exhibits a pronounced class imbalance (5:1 fake-to-real ratio).
+To address this challenge, we design a disjoint set-based multistage training strategy, where the fake class is divided into five non-overlapping subsets and trained sequentially along with real samples, using weight warm-starting across stages to retain learned knowledge.
+The final weighted ensemble employs probability-level late fusion to combine the outputs of the three backbones, leading to statistically significant improvements over individual models and conventional ensemble techniques.
+As we demonstrate in Section 3, the proposed model offers strong classification performance and maintains robustness under adversarial perturbations, confirming the effectiveness and reliability of CAE-Net in generalized deepfake image detection.
+The rest of the paper is organized as follows: Section 2 contains data preprocessing techniques, a description of our CAE-Net architecture, consisting of EfficientNet, DeiT, and ConvNeXt.
+Section 3 contains the classification results of multiple standard models on the dataset, an ablation study on the different ensemble variations, along with the implementation details and other evaluation metrics of our proposed approach.
 This also contains a comparison of our model’s generalization capabilities against multiple existing works and a defense against adversarial perturbations.
+Section  4 provides an overall view of the interpretation of the results, the significance of the findings, and the limitations of our study.
 Finally, Section 5 draws concluding remarks on the contributions presented in the paper.
@@
 One of the key contributions of our paper is to devise a disjoint set-based multistage training approach to address the class imbalance present in the given dataset.
+Traditional class-balancing techniques (oversampling, SMOTE, class weighting) present significant limitations for our dataset’s extreme imbalance.
+Random oversampling: We empirically validated that replicating 42,690 real images to match 219,470 fakes introduce severe overfitting during model training, as models see identical real samples multiple times across epochs, memorizing specific artifacts rather than learning generalizable features.
+Synthetic generation (SMOTE, GANs): Generating $\sim$177,000 synthetic real images risks introducing unrealistic artifacts or distribution shift.
+Moreover, since the task itself is deepfake detection, authentic real images must represent unmanipulated content.
+Synthetic augmentation could paradoxically teach models to accept fake-like patterns as real.
+Class weighting: While computationally efficient, loss reweighting alone does not prevent the model from being overwhelmed by the sheer volume of fake samples during training, leading to bias toward the majority class despite adjusted gradients.
+Our disjoint multistage approach addresses these issues by ensuring a balanced exposure where each stage trains on a 1:1 balanced subset (42,690 real + 43,894 fake images).
+All real images remain genuine and unmanipulated samples, preserving the true distribution of authentic content.
+This strategy effectively leverages our full dataset (262,160 images) without the artifacts, overfitting, or distribution shifts inherent to conventional balancing techniques applied to such extreme imbalances.
 Since the number of fake images is approximately five times the number of real images, we divided the fake images $\{\mathcal{F}\}$ into five disjoint subsets $\{\mathcal{F}_{1},\ldots,\mathcal{F}_{5}\}$.
-Each of the three aforementioned models, denoted as $M_{1},M_{2},M_{3}$, are trained in five stages with these disjoint subsets of fake images.
+Each of the three aforementioned models, denoted as $M_{1},M_{2},M_{3}$, is trained in five stages with these disjoint subsets of fake images.
 At each stage $i$, the model is trained using the entire real image set $\mathcal{R}$ and one subset of fake images $\mathcal{F}_{i}$, such that the training dataset of stage $i$ is given by:
 $\mathcal{D}_{i}=\mathcal{R}\cup\mathcal{F}_{i},\quad i\in\{1,2,3,4,5\}.$ (1)
-We emphasize that the same set of real images ${}\{\mathcal{R}\}$ are treated as one class during each stage.
+We emphasize that the same set of real images $\{\mathcal{R}\}$ is treated as one class during each stage.
 The training of each model is initialized with pre-trained weights $\theta_{0}$, obtained from an ImageNet [37]-trained version of the corresponding architecture.
@@
 This process continues until the model has been trained on all subsets of fake images, ensuring that the model learns from diverse variations of the fake images while preserving knowledge of real images.
+While training our models across stages, we maintained the hyperparameter settings stated in Table 8.
 2.2.2 Data Augmentation and Normalization
@@
 These augmentations enhanced the diversity of the training dataset, reduced overfitting, and improved the model’s ability to generalize across a wider range of deepfake detection scenarios.
-It should be noted that such augmentations were used only in training the DeiT architectures.
+It should be noted that such augmentations were used only in training the DeiT architecture.
+For EfficientNet, data augmentation didn’t enhance validation accuracy or generalization, likely due to its convolutional design already effectively captures the local features.
+In contrast, DeiT showed notable gains in validation accuracy with augmentation, as its attention-based architecture thrives on greater data variety to learn global patterns [38].
+Therefore, we applied augmentation only to DeiT in the final model.
 For the ConvNeXt, we employed wavelet transform-based feature extraction from the raw images, and data augmentation techniques would result in incorrect feature representation in the spectral domain.
@@
 The ability to decompose an image into different frequency sub-bands while preserving spatial relationships makes it particularly effective at identifying edges, textures, and subtle patterns of an image [40].
-Using 2D Wavelet transform, an image can be decomposed into approximate and detail coefficients [23].
+Using the 2D Wavelet transform, an image can be decomposed into approximate and detail coefficients [41].
 As such, we performed a single-level 2D Discrete Wavelet Transform (2D-DWT) using Haar wavelets to produce the approximate and detailed coefficients of the images and used the computed coefficients as features.
@@
 The Haar wavelet uses two filters:
-  • 1.
   • Low-pass filter ($\phi$): Coefficients $\left[\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right]$
-  • 2.
   • High-pass filter ($\psi$): Coefficients $\left[-\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}}\right]$
@@
     Column-wise filtering and downsampling:
-[…]
-[…]
-[…]
-[…]
+    $A$ $=\left(L_{\text{row}}\ast\phi\right)\downarrow_{2,\text{rows}}\qquad\text{(Approximation, LL)}$
+    $V$ $=\left(L_{\text{row}}\ast\psi\right)\downarrow_{2,\text{rows}}\qquad\text{(Vertical Details, LH)}$
+    $H$ $=\left(H_{\text{row}}\ast\phi\right)\downarrow_{2,\text{rows}}\qquad\text{(Horizontal Details, HL)}$
+    $D$ $=\left(H_{\text{row}}\ast\psi\right)\downarrow_{2,\text{rows}}\qquad\text{(Diagonal Details, HH)},$
 where $\ast$ indicates convolution; $\downarrow_{2,\text{cols}}$, and $\downarrow_{2,\text{rows}}$ indicate downsampling the columns and rows by 2, respectively; $A$ denotes low-frequency approximation coefficients; and $H,V,D$ denote high-frequency horizontal, vertical, and diagonal detail coefficients, respectively.
@@
 Figure 2: Real and Fake feature images after wavelet transform.
-The top left, top right, bottom left, and bottom right portions indicate the approximate, horizontal detail, vertical detail, and diagonal detail coefficients of the original image, respectively.
+The top left, top right, bottom left, and bottom right portions indicate the approximate horizontal detail, vertical detail, and diagonal detail coefficients of the original image, respectively.
 Different colormaps are applied to the horizontal, vertical, and diagonal detail coefficients, and their contrasts are increased for better visibility.
@@
 Squeeze-and-Excitation (SE) recalibrates channel importance by computing a channel-wise attention vector [43].
-The SE block acts as a lightweight attention mechanism to adaptively recalibrate each feature channel’s importance [37].
+The SE block acts as a lightweight attention mechanism to adaptively recalibrate each feature channel’s importance [44, 45].
 It is responsible for two main operations:
@@
 EfficientNet-B0 uses MBConv1 (expansion factor 1) to MBConv6 (expansion factor 6) variants, with kernel sizes of 3x3 or 5x5 depending on the stage.
-A brief summary of different layers, the output shapes and the number of parameters in the EfficientNet-B0 architecture is shown in Table 1.
+A summary of different layers, the output shapes, and the number of parameters in the EfficientNet-B0 architecture is shown in Table 1.
 Table 1: Summary of the EfficientNetB0 model architecture
@@
 2.4.2 Data Efficient Image Transformer (DeiT)
-Unlike conventional Recurrent Neural Networks (RNNs), transformers process all inputs simultaneously, relying on a self-attention mechanism.
+Unlike conventional Recurrent Neural Networks (RNNs), transformers process all inputs simultaneously, relying on a self-attention mechanism [47].
 This allows transformers to learn long-range dependencies and global contexts of data [48].
@@
 The summary number of parameters and the output shapes of layers are shown in Table 2.
-Table 2: Summary of the Data Efficient Image Transformer (DeiT) Model Architecture
+Table 2: Summary of the data-efficient image transformer (DeiT) model architecture
 Layer (type) Output Shape Param #
@@
 Non-Trainable Parameters 0
+Figure 3: Detail of the proposed CAE-Net architecture with weighted ensemble framework.
 2.4.3 ConvNeXt Model
@@
 Each ConvNeXt block contains a 7x7 depthwise convolution (one filter per input channel) that is applied to capture spatial features efficiently.
-Then layer normalization is applied before and after the convolution to stabilize training.
+Then, layer normalization is applied before and after the convolution to stabilize training.
 It normalizes features across the channel dimension for each spatial position.
@@
 A skip connection is used from the input to the block’s output, similar to ResNet’s residual connections [52].
-Between stages, ConvNeXt uses a downsampling layer with layer normalization and a 2x2 convolution with stride 2, inspired by Swin Transformer’s separate downsampling strategy, unlike ResNet’s within-block downsampling.
+Between stages, ConvNeXt uses a downsampling layer with layer normalization and a 2x2 convolution with a stride of 2, inspired by Swin Transformer’s separate downsampling strategy, unlike ResNet’s within-block downsampling.
 From the diverse family of ConvNeXt architectures, we specifically implemented the ConvNeXt-tiny model for its low compute and high accuracy on the dataset.
@@
 Conv2d + LayerNorm2d $[96,74,74]$ 4,896
-CNBlock (×$\times 3$) $[96,74,74]$ 237,600
+CNBlock ($\times 3$) $[96,74,74]$ 237,600
 LayerNorm2d + Conv2d $[192,37,37]$ 74,112
-CNBlock (×$\times 3$) $[192,37,37]$ 917,568
+CNBlock ($\times 3$) $[192,37,37]$ 917,568
 LayerNorm2d + Conv2d $[384,18,18]$ 295,680
-CNBlock (×$\times 9$) $[384,18,18]$ 10,813,824
+CNBlock ($\times 9$) $[384,18,18]$ 10,813,824
 LayerNorm2d + Conv2d $[768,9,9]$ 1,181,184
-CNBlock (×$\times 3$) $[768,9,9]$ 14,287,104
+CNBlock ($\times 3$) $[768,9,9]$ 14,287,104
 AdaptiveAvgPool2d + LayerNorm2d $[768,1,1]$ 1536
@@
 The associated weights for the EfficientNet, DeiT, and ConvNeXt with wavelet transform are 0.35, 0.34, and 0.31, respectively, based on our parameter search.
-The final architecture is shown in Figure 3.
-Figure 3: The proposed CAE-Net architecture
+Initially, each model was given equal weights (accuracy result for this selection of weights is stated in Table 9) and then manually tuned.
+To refine the fusion strategy, the weights were then manually tuned based on validation performance.
+Specifically, weight combinations were evaluated in 0.01 increments while satisfying the simplex constraint ($w_{1}+w_{2}+w_{3}=1$), and the configuration that achieved the highest accuracy on the SP Cup 2025 validation set was selected.
+This ensemble employs probability-level late fusion, where each model processes the input independently within its full architecture, preserving its native feature representation.
+The final architecture of CAE-Net is shown in Figure 3.
 2.5 Adversarial Perturbations
@@
 We systematically evaluated these models to test these hypotheses along with multiple other frequency domain feature extraction techniques, establishing the first performance benchmarks for the dataset.
-Table 4: Comparative Analysis of Different Model Variations
-Classifier Models Accuracy
-CNN models
-  VGG-19 [52] 62.25%
-  ResNet-50 [44] 69.00%
-  MobileNet-V2 78.12%
-  ResNet-34 [44] 80.32%
-  RegNet [53] 80.92%
-  Inception V3 [54] 84.38%
-  ResNet-34 with CBAM 85.19%
-  ResNet-34 with single SE 88.90%
-  DenseNet-121 [55] 90.01%
-  EfficientNet-B0 [21] 90.79%
-  EfficientNet-B4 [21] 87.27%
-  EfficientNet-B5 [21] 88.31%
-  ResNeXt [56] 85.19%
-  ConvNeXt [43] 88.18%
-Frequency Domain Experimentations
-  DCT with ResNet-34 71.48%
-  ResNet-34 with Wavelet 80.44%
-  EfficientNetB0 with Wavelet 82.48%
-  Xception-Net [57] with Wavelet 87.76%
-  ConvNeXt with Wavelet 89.49%
-Study on Transformers
-  ViT-B [58] 78.12%
-  Swin Transformer [45] 86.52%
-  MobileViT [59] 87.04%
-  DeiT [39] 89.32%
-Ensemble Results
-  Maj. Voting (EfficientNet, DeiT, ConvNeXt) 94.24%
-  Proposed Weighted Ensem. (EfficientNet, DeiT, ConvNeXt) 94.63%
-Hyperparameters.
-For the training of our models, we have used the ADAM optimizer and binary categorical cross-entropy as our loss function.
-Using the validation set to observe the accuracy of the proposed architecture, each image took approximately 0.03385416s on a Tesla P100 GPU.
-The hyperparameters used for training are presented in Table 5.
+We note that all the results noted for different benchmark models in Table 456 are done with a fixed random seed of 42.
+Table 4: Comparative analysis of different CNN model variations to capture local artifacts
+Classifier Models Accuracy # of params.
+VGG-19 [7] 62.25% 144 M
+ResNet-50 [52] 69.00% 25.6 M
+MobileNet-V2 [42] 78.12% 3.5 M
+ResNet-34 [52] 80.32% 21.3 M
+RegNet-400mF [60] 80.92% 5.2 M
+Inception V3 [61] 84.38% 26 M
+ResNet-34 with CBAM [^1] Convolutional Block Attention Module 85.19% 24.5 M
+ResNet-34 with SE [^2] Squeeze and Excitation 88.90% 21.3 M
+DenseNet-121 [62] 90.01% 7.98 M
+EfficientNet-B0 [26] 90.79% 3.96 M
+EfficientNet-B4 [26] 87.27% 19.3 M
+EfficientNet-B5 [26] 88.31% 30.4 M
+ResNeXt-50 [4] 85.19% 25 M
+ConvNeXt [28] 88.18% 27.8 M
+The performance and size comparison of different benchmark CNN models are shown in Table 4.
+VGG-19 is the largest model by a significant margin, but has the lowest accuracy.
+The use of attention mechanisms like CBAM and SE blocks in the ResNet-34 variants appears beneficial, as the 88.90% accuracy of ResNet-34 with SE is substantially higher than the baseline ResNet-34 (80.32%).
+This led to our choice of EfficientNet, which has in-built SE blocks and is less compute-heavy compared to other CNNs.
+Among different variants of EfficientNets, EfficientNet-B0 has the highest performance with the lowest number of parameters, making it an optimum choice for a backbone model capturing local artifacts.
+DenseNet and ConvNeXt show high performance as well, making them good candidates for the frequency domain experimentations.
+Table 5: Comparative analysis of different transformer model variations to capture global artifacts
+Classifier Models Accuracy # of params.
+ViT-B [8] 78.12% 86.6 M
+Swin Transformer [53] 86.52% 28.3 M
+MobileViT [63] 87.04% 5.6 M
+DeiT-B [27] 89.32% 85.6 M
+As shown in Table 5, the ViT-B and DeiT-B models are the largest, both with parameter counts around 86 M.
+Despite being similar in size, DeiT-B significantly outperforms ViT-B in accuracy, suggesting the training strategies used in DeiT are highly effective.
+Among the transformers, MobileViT is the most parameter-efficient model using only 5.6 M parameters.
+However, since achieving the highest accuracy is of the utmost importance in sensitive issues like deepfake detection, we take DeiT-B as a backbone model to capture global artifacts.
+Table 6: Comparative analysis of different CNN model variations to capture frequency domain artifacts
+Classifier Models Accuracy # of params.
+ResNet-34 with FFT [64] 66.55% 21.3 M
+ResNet-34 with DCT [11] 71.48% 21.3 M
+ResNet-34 with Wavelet [39] 80.44% 21.3 M
+EfficientNetB0 [26] with Wavelet 82.48% 3.96 M
+RegNet-Y [60] with Wavelet 86.04% 10.31 M
+Xception-Net [5] with Wavelet 87.76% 20.8 M
+ConvNeXt [28] with Wavelet 89.49% 27.81 M
+Additionally, we conducted a comparative analysis of different frequency domain feature extraction techniques, such as DCT, Fast Fourier Transform (FFT), and 2D Discrete Wavelet Transform (DWT), as preprocessing steps with benchmark CNN models and the results are shown in Table 6.
+DCT and FFT with ResNet-34 yield very low accuracy among the specialized models.
+In contrast, ResNet-34 with Wavelet significantly improves performance to 80.44%, suggesting that the Wavelet transform is generally more effective than the DCT for this specific classification task.
+Therefore, we experimented with other high-performing CNN models with the wavelet transform as a pre-processing technique.
+The ConvNeXt with Wavelet model achieves the highest accuracy at 89.49%.
+Therefore, we take this model as a backbone for capturing frequency domain artifacts specifically.
+3.1 Choice of Hyperparameters
+Due to computational resource constraints, a comprehensive hyperparameter tuning using advanced optimization algorithms was not feasible.
+Instead, we conducted a discrete evaluation of key hyperparameters to identify optimal settings for the first stage of our five-stage training process, which spanned five epochs.
+We assessed batch sizes of 16, 32, and 64 (each with a fixed learning rate of $1\times 10^{-4}$) and learning rates of $\times 10^{-3}, 1×1$$\times 10^{-4}$, and $1\times 10^{-5}$ (each with a fixed batch size of 32).
+This analysis focused on recording the validation accuracy of EfficientNet, DeiT, and ConvNeXt for each configuration.
+The results are presented in Table 7, which guided the selection of the most effective settings, balancing performance and resource efficiency.
+We acknowledge this as a limitation and plan to explore more extensive tuning in future work with adequate resources.
+Table 7: Accuracy results for different hyperparameter configurations
+Parameter Accuracy
+EfficientNet DeiT ConvNeXt
+Batch Size:
+         16 91.40% 89.78% 87.37%
+         32 90.79% 89.32% 89.49%
+         64 87.34% 84.86% 91.41%
+Learning Rate:
+[…]
+[…]
+[…]
+As shown in Table 7, EfficientNet and DeiT achieved the highest accuracy for a batch size of 16, while ConvNeXt performed the best with a batch size of 64.
+For learning rate selection, the CNN models performed well at $1\times 10^{-4}$, whereas DeiT achieved its best result at $1\times 10^{-5}$.
+Considering these observations, we selected a learning rate of $1\times 10^{-4}$ with a batch size of 32 for ensemble training, providing a balanced configuration that supports stable convergence across all three models and results in improved ensemble accuracy.
+For training, we used the ADAM optimizer with binary categorical cross-entropy as the loss function.
+All experiments were conducted on an NVIDIA Tesla P100 GPU.
+The final hyperparameters used for model training are summarized in Table 8.
+We have implemented similar hyperparameter settings for each stage in our five-stage training approach.
+Each stage was trained with an equal number of epochs to ensure consistency, and the images of both classes went under augmentation only while training our DeiT architecture.
 Table 8: Hyperparameters for model training
@@
 Batch size 32
-Number of Epochs in each session
-   - EfficientNet 5
-   - DeiT 5
-   - ConvNeXt 4
+Number of Epochs in Each Stage
+     EfficientNet 5
+     DeiT 5
+     ConvNeXt 4
 Learning rate
-   - Learning rate scheduler ReducedLROnPlateau
-   - Initial learning rate 0.0001
-   - Reduction factor 0.1
-   - Patience 3
+     Learning rate scheduler ReducedLROnPlateau
+     Initial learning rate 0.0001
+     Reduction factor 0.1
+     Patience 3
+3.2 Ablation study
+Table 9: Ablation study on ensemble for best performing models
+EfficientNet DeiT ConvNeXt Ensemble # of params FLOPs Inference Throughput Accuracy
+(M) (G) time (ms) (image/s) (%)
+3.96 0.386 24.42 1310.26 86.11 $\pm$ 3.22
+85.64 16.863 178.57 179.2 89.16 $\pm$ 1.26
+27.81 7.521 128.81 248.43 91.24 $\pm$ 1.22
+equal weights 89.61 17.247 202.95 157.67 92.59 $\pm$ 1.87
+equal weights 31.77 7.907 153.11 209.0 93.14 $\pm$ 1.43
+equal weights 113.46 24.384 307.64 104.02 91.87 $\pm$ 0.78
+equal weights 117.426 24.768 332.08 96.42 94.16 $\pm$ 0.61
+optimized weights 117.426 24.768 332.08 96.42 94.46 $\pm$ 0.58
+majority voting 117.426 24.768 332.08 96.42 94.09 $\pm$ 0.53
+Ablation study on ensemble configurations.
+To better understand the individual contributions of each sub-module in the proposed ensemble, we conducted a detailed ablation study using the three backbone models, namely the EfficientNet, DeiT, and ConvNeXt, across various combinations.
+The analysis aimed to examine not only the detection performance but also the practical deployability of each configuration by reporting the number of parameters, floating point operations (FLOPs), inference time, and throughput (measured with a batch size of 32).
+Three ensemble strategies were evaluated: (i) equal weights, where all models contributed equally; (ii) optimized weights, where coefficients were manually tuned on the fixed validation set; and (iii) majority voting, which relied on discrete decision fusion.
+Each setting was trained and tested five times with random initializations, and the average accuracy with standard deviation is summarized in Table 9.
+From the results, the individual models showed competitive but varying performance, with ConvNeXt achieving the highest single-model accuracy (91.24%) among the three.
+Combining multiple models consistently improved results, demonstrating the complementary strengths of spatial (EfficientNet), global (DeiT), and frequency-domain (ConvNeXt + wavelet) features.
+The optimized weighted ensemble achieved the best overall accuracy (94.46%), outperforming both equal weighting and majority voting approaches.
+In addition to performance, we included deployment-oriented metrics, such as parameter count, FLOPs, throughput, and inference time, to provide a clearer view of the trade-off between accuracy and computational cost.
+These added measurements help assess the real-world feasibility of the model, showing that the proposed ensemble maintains strong efficiency while offering significant accuracy gains over its individual components.
 Result Analysis.
@@
 While calculating these metrics, real images were considered as class 0 and fake images as class 1.
-The performance of our model on these metrics is presented in Table 6 and the confusion matrix and the ROC curve are shown in Figure 4.
+The performance of our model on these metrics is presented in Table 10, and the confusion matrix and the ROC curve are shown in Figure 4.
+The results shown here are for a fixed seed of 42.
 It is evident from the evaluation indices that our proposed weighted ensemble architecture performs very well on the competition dataset, which underlines the generalization capability of the proposed approach.
@@
 Evaluation Metrics Performance
-Accuracy 94.63%
-Precision 0.937
-Recall 0.956
-F1 score 0.9464
-Area under ROC 0.9737
-Equal Error Rate 4.72%
-Average Precision 0.9771
-Figure 4: (a) Confusion matrix and (b) ROC curve for validation set
+Accuracy 94.46%
+Precision 93.70%
+Recall 95.60%
+F1 score 94.64%
+Area under ROC 97.60%
+Equal Error Rate 5.00%
+Average Precision 97.71%
+Figure 4: (a) Confusion matrix and (b) ROC curve of the proposed CAE-Net evaluated on the SP Cup 2025 validation set.
+3.3 Interpretability analysis of the proposed network
 To investigate the performance of our ensemble model, we intend to visualize i) the regions of the images that the three models focus on for classification, and ii) how the three models distinguish real and fake images before and after training.
-To that end, we note that the Grad-CAM visualizations show how different models focus on different facial regions of the same image for deepfake detection.
+To that end, we note that the Grad-CAM visualizations show how different models focus on different facial regions for deepfake detection.
 In Figure 5, it can be seen that EfficientNet focuses on localized facial features, whereas DeiT displays diffuse attention patterns indicating a focus on global features.
@@
 Additionally, we present the Grad-CAM view of some of the wrong predictions by the three models in Figure 5, showing the focus of irrelevant regions, leading to misclassifications.
-(a) EfficientNet
-(b) DeiT
-(c) ConvNeXt
-Figure 5: Importance map shown by Grad-CAM for different models
-(a) EfficientNet
-(b) DeiT
-(c) ConvNeXt
-Figure 6: Grad-CAM view of a few failed classifications of our proposed model
-(a) EfficientNet (Untrained)
-(b) DeiT (Untrained)
-(c) ConvNeXt (Untrained)
-(d) EfficientNet (Trained)
-(e) DeiT (Trained)
-(f) ConvNeXt (Trained)
-Figure 7: t-SNE plot of three models before and after training for randomly selected 4000 real and fake images using perplexity = 40
+Figure 5: Importance map shown by Grad-CAM for different models with correct and wrong predictions.
+The green boxes indicate the grad-CAM view for correct predictions, and the red boxes indicate the wrong ones.
 Next, we note that high-dimensional data can be visualized by projecting onto a two-dimensional space using the non-linear dimensionality reduction technique known as t-SNE (t-Distributed Stochastic Neighbor Embedding) [65].
-By maintaining the local neighborhoods of data points, it shows the differences and similarities of data points in the learned feature space.
-As such, the t-SNE visualizations (Fig.
-7) offer valuable insights into the effectiveness of our models.
+Maintaining the local neighborhoods of data points shows the differences and similarities of data points in the learned feature space.
+As such, the t-SNE visualizations (Figure 6) offer valuable insights into the effectiveness of our models.
 For both real and fake images, we observe cohesive clusters for all three models, despite the images being produced using various deepfake generation methods.
 This highlights the generalization capability of our proposed ensemble model.
-After training, all three backbone models, i.e., EfficientNet, DeiT, and ConvNeXt, produced well-separated embeddings for real and fake images, as illustrated in Fig.
-7(d), 7(e), and 7(f).
+After training, all three backbone models, i.e., EfficientNet, DeiT, and ConvNeXt, produced well-separated embeddings for real and fake images, as illustrated in Figure 6(d), (e), (f).
+Figure 6: t-SNE plot of three models before and after training for randomly selected 4000 real and fake images using perplexity = 40.
+(a) EfficientNet (Untrained), (b) DeiT (Untrained), (c) ConvNeXt (Untrained), (d) EfficientNet (Trained), (e) DeiT (Trained), and (f) ConvNeXt (Trained).
+To quantitatively assess the cluster separability of real and fake embeddings in the t-SNE space, four widely used metrics were considered: Calinski–Harabasz (CH) Index [66], Davies–Bouldin (DB) Index [67], Inter-Centroid Euclidean Distance, and Inter-Centroid Mahalanobis Distance [68].
+The CH Index measures the ratio of between-cluster dispersion to within-cluster compactness—higher values indicate better separation.
+The DB Index reflects average cluster similarity, where lower values denote reduced overlap.
+The Inter-Centroid Euclidean Distance quantifies how far apart the mean feature representations of two classes are, while the Mahalanobis Distance accounts for covariance, offering a more distribution-aware measure of separation.
+Table 11: Cluster separability metrics (t-SNE space) for different models before and after training.
+Model Calinski–Harabasz $\uparrow$ Davies–Bouldin $\downarrow$ Inter-Centroid (Euclidean) $\uparrow$ Inter-Centroid (Mahalanobis) $\uparrow$
+EfficientNet (Untrained) 10.9841 16.3122 1.0779 1.9434
+EfficientNet (Trained) 358.1060 2.8366 2.5301 1.9628
+DeiT (Untrained) 2.5970 36.9521 1.9842 0.0643
+DeiT (Trained) 6799.6961 0.7083 68.3502 1.7020
+ConvNeXt (Untrained) 4.5949 27.3898 2.9342 0.1013
+ConvNeXt (Trained) 4530.9877 0.8663 61.1882 1.7710
+As shown in Table 11, all three models demonstrate substantial improvements in separability after fine-tuning.
+For example, the Calinski–Harabasz score rises from 10.98 to 358.10 for EfficientNet, from 2.60 to 6799.70 for DeiT, and from 4.59 to 4530.99 for ConvNeXt, indicating a remarkable increase in inter-class distinction.
+Correspondingly, the Davies–Bouldin index decreases sharply across all models, confirming tighter and more compact clusters.
+Both Euclidean and Mahalanobis distances between class centroids also increase, showing that fine-tuning pushes the real and fake distributions farther apart in feature space.
+Together, these results provide strong quantitative evidence that the proposed training strategy significantly enhances feature discriminability, leading to more distinct and separable clusters in the t-SNE representation.
+3.4 Statistical analysis
+To further validate the effectiveness of the proposed ensemble model, we conducted a pairwise statistical comparison using the McNemar test.
+This non-parametric test assesses whether two classifiers differ significantly in their performance on paired nominal data, focusing specifically on instances where the models disagree.
+The results of this analysis are summarized in Table 12.
+Among the individual models, EfficientNet shows slightly better performance than ConvNeXt ($p=0.047$), with 40 more correct predictions in disagreement cases.
+However, this difference does not meet the Bonferroni-corrected significance threshold ($\alpha=0.008$).
+A statistically significant difference is observed between EfficientNet and DeiT ($p=0.003$), where EfficientNet outperforms DeiT by 58 correct predictions in disagreement cases.
+In contrast, ConvNeXt and DeiT show no significant difference ($p=0.389$), suggesting comparable decision behavior.
+These observations indicate that the three backbone models capture distinct yet partially overlapping feature representations of deepfake artifacts, reflected in their areas of disagreement.
+Table 12: Results of McNemar statistical test comparing pairwise classifier performance
+Model 1 Model 2 $n_{11}$ $n_{10}$ $n_{01}$ $n_{00}$ $\chi^{2}$ p-value Significant
+(both correct) (M1 correct) (M2 correct) (both wrong) ($\alpha=0.05$)
+EfficientNet ConvNeXt 2576 213 173 110 3.94 0.047 Yes
+EfficientNet DeiT 2572 217 159 124 8.641 0.003 Yes
+ConvNeXt DeiT 2545 204 186 137 0.741 0.389 No
+EfficientNet CAE-Net 2752 37 145 138 62.907 0.000 Yes
+ConvNeXt CAE-Net 2720 29 177 146 104.898 0.000 Yes
+DeiT CAE-Net 2701 30 196 145 120.465 0.000 Yes
+The ensemble model (CAE-Net), however, demonstrates a clear and statistically significant improvement over all three individual models, with $p$-values approaching zero even after Bonferroni correction.
+Specifically, the ensemble achieves 108, 148, and 166 additional correct classifications compared to EfficientNet, ConvNeXt, and DeiT, respectively, in the disagreement cases where differential performance matters most.
+Moreover, the relatively low number of disagreements between the ensemble and base models (182–226 cases) suggests that the ensemble rarely errs when base models succeed, but frequently corrects their mistakes.
+Overall, these results provide strong statistical evidence that the ensemble’s weighted fusion mechanism effectively leverages the complementary strengths of its constituent architectures.
+The ensemble not only achieves higher accuracy but also exhibits more consistent and reliable decision-making, confirming that its performance gains are statistically robust and unlikely due to chance.
+3.5 Comparison with the existing works
 There exist very few works on deepfake detection that perform well on the combination of such a diverse set of datasets.
@@
 Note that the AUC score presented here is the mean score of these models on the corresponding datasets (we refer the reader to Figure 8 of [30]).
-We want to emphasize that existing works often demonstrate good performance on only a few datasets, but the overall performance on multiple datasets is very poor for most architectures.
+We want to emphasize that existing works often demonstrate good performance on only a few datasets, but the overall performance on multiple datasets is poor for most architectures.
 Table 13: Comparative analysis of average AUC score of different models
-[…]
-[…]
+  Detection Method Datasets AUC Score Std. Dev.
+  Capsule [69] UADFV [34], DF-TIMIT [70], FaceForensics++ [31], 69.4% 13.84%
 DFD [32], DFDC [36], Celeb-DF [30]
@@
 FaceShifter [33], DFDC Preview [35], FaceForensics++ [31]
-Finally, we look into the effectiveness of our ensemble method through a simple adversarial perturbation attack, such as Fast Gradien Sign Method (FGSM).
+We compared our proposed model with existing ensemble and frequency–spatial hybrid architectures on the DF-Wild Cup dataset to evaluate both classification performance and model complexity.
+As shown in Table 14, CAE-Net achieves superior performance, attaining 94.46% accuracy with 117.4M parameters, outperforming previous methods, such as ResNeXt–Xception, DeepfakeStack, and CoAt-Net ensembles.
+Unlike prior approaches that primarily focus on either spatial or hybrid CNN–Transformer integration without frequency-domain enhancement, CAE-Net incorporates wavelet-based preprocessing to extract complementary frequency cues while maintaining architectural diversity among its sub-models.
+This balanced design allows the ensemble to learn richer representations of deepfake artifacts and to generalize more effectively across diverse manipulation techniques.
+Consequently, CAE-Net achieves a better trade-off between accuracy and computational cost, demonstrating the effectiveness of its ensemble logic compared to existing state-of-the-art models.
+Table 14: Comparative analysis of our work with prior ensembles and frequency-spatial hybrid networks
+Classifiers # of params. Accuracy
+Freq-Net [73] 1.85 M 67.35%
+FSBI [21] 23.34 M 82.00%
+CoAt-Net Ensemble [74] 699.72 M 84.22%
+SFA-Net [75] 25.04 M 90.07%
+ResNeXt & Xception Ensemble [3] 43.8 M 91.21%
+DeepfakeStack [76] 135.55 M 92.12%
+Proposed CAE-Net 117.4 M 94.46%
+3.6 Robustness analysis against adversarial attacks
+Finally, we look into the effectiveness of our ensemble method through a simple adversarial perturbation attack, such as Fast Gradient Sign Method (FGSM).
 As mentioned before, FGSM can fool models by adding simple perturbation noise based on the gradient sign that intends to maximize the loss function of the detection model instead of minimizing it.
 We present the accuracy scores of the separate detection model as well as the ensemble models for different values of $\epsilon$ in Table 15.
-Table 8: Comparison of Our Proposed Architectures under Adversarial FGSM Attacks
-Model Accuracy
-$\epsilon=0.01$ $\epsilon=0.03$ $\epsilon=0.05$
-EfficientNetB0 38.54% 46.48% 53.94%
-Data Efficient Image Transformer 42.76% 21.94% 16.73%
-ConvNeXt with wavelet 6.45% 6.51% 15.40%
-Weighted Ensem.(EfficientNet, DeiT, ConvNeXt) 12.04% 13.48% 16.05%
-Majority Voting (EfficientNet, DeiT, ConvNeXt) 25.52% 20.41% 19.76%
+Table 15: Comparison of our proposed architectures under FGSM attacks
+Model Accuracy (before adversarial training) Accuracy (after adversarial training)
+$\epsilon=0.005$ $\epsilon=0.01$ $\epsilon=0.03$ $\epsilon=0.05$ $\epsilon=0.00$ $\epsilon=0.005$ $\epsilon=0.01$ $\epsilon=0.03$ $\epsilon=0.05$
+EfficientNetB0 42.77% 38.54% 46.48% 53.94% 93.07% 56.90% 50.39% 45.21% 50.26%
+DeiT 49.61% 42.76% 21.94% 16.73% 89.29% 65.10% 37.14% 5.57% 3.74%
+ConvNeXt with wavelet 16.08% 6.45% 6.51% 15.40% 92.25% 98.40% 96.45% 63.28% 51.63%
+Weighted Ensem.(CAE-Net) 20.64% 12.04% 13.48% 16.05% 94.43% 87.27% 69.92% 14.84% 20.25%
+Majority Voting (CAE-Net) 34.86% 25.52% 20.41% 19.76% 93.85% 82.58% 64.29% 25.03% 35.52%
 The results show an increased robustness of the EfficientNet architecture, which depicts increased accuracy with increasing perturbation compared to the other two architectures.
-DeiT shows a decrease in accuracy with rising perturbation, and ConvNeXt accuracy drops significantly even for very low perturbation added to the image.
+DeiT shows a decrease in accuracy with rising perturbation, and ConvNeXt accuracy drops significantly even for very low perturbations added to the image, which is likely caused by the modification of frequency domain artifacts by even small attacks, which disrupts its feature extraction.
 Though a weighted ensemble model is appropriate for achieving high accuracy on the generalized dataset, we argue that it is highly vulnerable to adversarial attacks.
 If any model performs poorly on adversarial data, then the performance of the entire ensemble drops significantly.
-However, in the majority voting ensemble, the performance does not reduce as much, since it requires the majority of models to perform poorly, implying increased robustness against adversarial attacks for applications, such as deepfake detection.
+However, in the majority voting ensemble, the performance does not reduce as much, since it requires the majority of models to perform poorly, implying moderate robustness against adversarial attacks for applications, such as deepfake detection.
+To improve the robustness of our models against adversarial attacks, we integrated adversarial training using pre-trained weights optimized for clean data.
+This involved training with adversarial samples generated using a small perturbation ($\epsilon=0.005$) over six epochs, employing a reduced learning rate of 1e-5 and a batch size of 64.
+We adopted a mixed batch strategy, comprising 32 perturbed images and 32 unperturbed images per batch, to balance exposure to both conditions [57].
+Notably, this approach yielded significant enhancements, with models achieving improved accuracy on the clean validation set and demonstrating superior resilience to smaller FGSM perturbations ($\epsilon=0.005,0.01$) without adversarial training.
+The weighted ensemble and majority voting also show enhanced resilience to the FGSM attacks, with accuracy shifting from 20.64% without adversarial training to 87.27% (see Table 15) after adversarial training for $\epsilon=0.005$.
+The increase in overall accuracy is maintained for higher perturbations as well.
+Again, majority voting outperforms weighted ensemble in the case of higher perturbations, which is expected.
+Further training with incrementally larger perturbations and with more sophisticated adversarial images (such as, projected gradient descent) may enhance performance against higher perturbation levels, a direction we intend to explore in future work.
+4 Discussion
+This study presents CAE-Net, an ensemble architecture combining convolutional and attention-based models with spatial and frequency domain features for generalized deepfake detection.
+In this section, we discuss our design decisions, analyze the deployment implications of our results, acknowledge the limitations of our work, and outline future research directions.
+The first major challenge of our study was addressing the 5:1 class imbalance in our dataset.
+Initially, we attempted conventional techniques like random oversampling and weighted loss functions.
+However, these techniques presented critical limitations, as stated in Section 2.2.1.
+Our method ensures balanced 1:1 exposure at each of the five training stages, with each stage using all 42,690 real images paired with a disjoint subset of 43,894 fake images.
+Through this method, the authenticity of real images is preserved, and knowledge accumulation is continued via warm starting the next stages.
+Recent estimates suggest AI-generated content now exceeds human-created content in volume online [77].
+As deepfake generation becomes more accessible, detection datasets will increasingly exhibit class imbalance, with synthetic samples vastly outnumbering authentic ones.
+Our disjoint multistage training method provides a scalable solution for such scenarios without requiring proportional increases in authentic training data.
+We position CAE-Net for deployment scenarios where accuracy is very important and computational resources are available, such as centralized content moderation systems or forensic laboratories.
+For resource-constrained environments, such as edge devices, the individual EfficientNet model provides a viable alternative with only 3.96 M parameters and 386 M FLOPs, as shown in Table 9.
+For scenarios where resource constraints are more relaxed, the ensemble of EfficientNet and ConvNeXt is also a good option with a 1.5s inference time and 31.77 M parameters with a high accuracy of 93.14%.
+Note that the ensemble of EfficientNet and ConvNeXt provides detection accuracy comparable to CAE-Net with less than one-fourth the parameters and half the inference time.
+However, we emphasize here the 1.32% accuracy improvement of CAE-Net is equivalent to correctly identifying 13,200 additional deepfakes per million processed images in content moderation systems.
+Given the serious consequences of deepfake spread, including financial fraud, electoral interference, and non-consensual imagery, the requirement for high accuracy justifies our ensemble approach’s computational expense in critical applications.
+Our evaluation under FGSM attacks revealed that CAE-Net’s accuracy drops significantly under perturbations for the weighted ensemble, while majority voting performs slightly better under the same attack.
+This vulnerability stems from our probability-level fusion mechanism: when adversarial perturbations cause even one model’s output probabilities to shift significantly, the weighted sum propagates this error proportionally.
+However, the binary decision mechanism of majority voting scheme provides a bounded impact where a single compromised model cannot override the other two.
+This represents a fundamental trade-off between clean-data optimization (weighted ensemble) and adversarial resilience (majority voting).
+However, upon adversarial training, the models show better robustness against such attacks, outlining the importance of such enhancement methods for deepfake classifiers.
+For deployments in adversarially exposed environments such as public-facing platforms, we recommend majority voting.
+For accuracy-critical applications with lower adversarial threat, the weighted ensemble remains preferable.
+CAE-Net employs a probability-level fusion rather than intermediate feature-level fusion.
+Each model (EfficientNet, ConvNeXt, DeiT) processes inputs independently through its complete architecture, maintaining native feature representations without cross-model feature map alignment or shared layers.
+This design choice offers key advantages: architectural independence allows each model to be trained and optimized separately, computational efficiency is obtained through parallel inference, and the modularity of late fusion enables straightforward addition or removal of ensemble members.
+Fusion occurs only at the final layer through weighted averaging of class probabilities, making the ensemble mechanism interpretable and the contribution of each model directly quantifiable.
+Limitations and Future Directions.
+Several limitations of our work require discussion.
+First, the dataset provided by the IEEE SPCup 2025 organizers had a fixed validation set.
+While defining the weights of the ensemble, we could not implement cross-validation due to this fixed split, which limits the robustness of our weight selection.
+Second, hyperparameter optimization was performed manually due to resource constraints rather than through systematic grid search or automated methods.
+Third, while testing the benchmark model performances described in Table 456, we used a fixed random seed of 42 for reproducibility and fair comparison.
+While this ensures consistent results for that specific initialization, it does not guarantee that our observed improvements generalize across different random initializations.
+The Grad-CAM visualizations indicate that the models sometimes concentrate on irrelevant areas, such as the background or non-facial features, instead of the artifact-prone facial regions essential for accurate classification.
+This misdirected focus contributes to misclassifications.
+To address this issue, incorporation of attention regularization techniques can work to guide the models toward critical facial areas, prioritizing zones with significant artifacts.
+Additionally, our evaluation does not assess performance on compressed or degraded images, which are common in real-world social media contexts.
+The robustness of frequency-domain features to such compression artifacts requires further investigation.
+Our ablation studies focused on architectural combinations but did not explore alternative fusion strategies beyond weighted averaging and majority voting, such as learned meta-classifiers, attention-based fusion, or stacking.
+Future research directions include developing efficient model compression techniques, such as knowledge distillation, quantization, and pruning, to enable deployment on resource-constrained devices while maintaining high accuracy, and testing generalization to emerging deepfake generation methods, including diffusion models and neural radiance fields (NeRFs) [78].
 5 Conclusion
@@
 We emphasize that detecting deepfake images that are generated with a range of generative AI approaches is a challenging task.
-To that end, we have addressed a significant class imbalance through disjoint training set-based multistage training, and proposed a weighted ensemble approach that takes advantage of the strengths of an EfficientNet model, a Data Efficient Image Transformer model, and a ConvNext model for the accurate classification of deepfake images.
+To that end, we have addressed a significant class imbalance through disjoint training set-based multistage training and proposed a novel weighted ensemble approach that takes advantage of the strengths of an EfficientNet model, a Data Efficient Image Transformer model, and a ConvNext model for the accurate classification of deepfake images.
 The EfficientNet-B0 model enabled extracting local features from an image, while the DeiT model specialized in capturing global features.
 Additionally, ConvNeXt, combined with the wavelet transform, effectively identified frequency-dependent artifacts.
+Statistical analysis shows that the CAE-Net significantly outperforms the constituent backbone models.
 Through t-SNE plots, we demonstrated that our model is capable of separating real and fake images into separable clusters.
@@
 Besides, we further analyzed the effectiveness of the proposed models by testing them against FGSM attacks and found the optimum ensemble mechanism to be majority voting compared to the weighted ensemble with a little trade-off in accuracy.
-An interesting future work could be incorporating lightweight architectures, reducing the computational overhead, and increasing the model’s adaptability to newer and sophisticated deepfake generation methods.
+An interesting future direction could be the incorporation of lightweight architectures, which would reduce computational overhead and increase the model’s adaptability to newer and more sophisticated deepfake generation methods.